### v1

# Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

Abstract
Label smoothing (LS) is a popular regularisation method for training deep neural network classifiers due to its effectiveness in improving test accuracy and its simplicity in implementation. “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting. In this work, we reveal that LS negatively affects selective classification (SC) – where the aim is to reject misclassifications using a model’s predictive uncertainty. We first demonstrate empirically across a range of tasks and architectures that LS leads to a consistent degradation in SC. We then explain this by analysing logit-level gradients, showing that LS exacerbates overconfidence and underconfidence by regularising the max logit more when the probability of error is low, and less when the probability of error is high. This elucidates previously reported experimental results where strong classifiers underperform in SC. We then demonstrate the empirical effectiveness of logit normalisation for recovering lost SC performance caused by LS. Furthermore, based on our gradient analysis, we explain why such normalisation is effective. We will release our code shortly.
Keywords:
Uncertainty Estimation Selective Classification Label Smoothing
1 Introduction
Label smoothing (LS) [70] is a common regularisation technique used to improve classification accuracy in supervised deep learning. When training with cross-entropy minimisation, one-hot labels are linearly combined with a uniform distribution over classes, redistributing the probability mass and “smoothing” the “hard” targets. This discourages overfitting on the training labels. Due to its simplicity and empirical effectiveness, label smoothing features in many recent training recipes [75, 71, 28, 74, 52, 51, 72, 73, 53], being particularly popular on the ImageNet-1k [63] image classification benchmark.
Within the domain of uncertainty estimation, label smoothing is well explored in the context of model calibration [59, 10, 58, 47], where the aim is to align a model’s output probabilities with its empirical accuracy. The pairing is intuitive as label smoothing encourages models to be less confident, and models are typically more confident than they are accurate [25]. On the other hand, there is a distinct lack of research investigating the effects of label smoothing in the context of Selective Classification. Selective Classification (SC) [30, 22, 81, 36] is a problem setting where, in addition to the primary classification task, a binary rejection decision is made based on the predictive uncertainty estimated by the model, i.e. reject if uncertain. The aim is to reduce the number of errors served by the classifier by pre-emptively rejecting potential misclassifications. It is well motivated by applications where safety and reliability are important due to the high cost of failure, such as autonomous driving [37] and medical diagnosis [1, 42].
[…]
Recent large-scale empirical evaluations of pre-trained models [20, 4] have shown that many strong classifiers have surprisingly poor SC ability. That is to say, in spite of having very low error rates, they are poor at rejecting their own misclassifications. Unfortunately, by evaluating pre-trained models from open-source repositories, the authors find it difficult to disentangle any potential causes of this behaviour, as model training recipes vary wildly from checkpoint to checkpoint. In this work, we reveal that label smoothing is one such potential cause and present the following key contributions:
  1. 1.
    We show empirically, across a range of network architectures and vision tasks, that training with LS consistently leads to degraded SC performance, even if it may improve accuracy (see Fig. 1). As LS can be found in the training recipes of many of the models evaluated in [20, 4], this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
  2. 2.
    We are able to explain the above behaviour by analysing the logit-level gradients of the LS loss. We show that the amount LS regularises the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing as $P_{\text{error}}$ decreases. This exacerbates overconfidence and underconfidence, degrading SC.
  3. 3.
    Finally, we demonstrate that logit normalisation proposed by Cattelan and Silva [4] is highly empirically effective in combating specifically the degradation caused by LS. Moreover, we are able to explain the effectiveness of logit normalisation through the lens of our gradient-based analysis, showing that normalisation compensates for the imbalanced regularisation by increasing uncertainty as the max logit increases. [^1] Although they demonstrate its empirical effectiveness, Cattelan and Silva [4] do not make a connection to label smoothing, and provide limited insight as to why their approaches are effective.
2 Preliminaries
Consider a $K$-class neural network classifier with parameters $\boldsymbol{\theta}$ that models the conditional distribution $P(y|\boldsymbol{x};\boldsymbol{\theta})$ of labels $y\in\mathcal{Y}=\{\omega_{k}\}_{k=1}^{K}$ given inputs $\boldsymbol{x}\in\mathcal{X}=\mathbb{R}^{D}$. Typically the network has a categorical softmax output $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\in[0,1]^{K}$,
$P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})=\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=\exp v_{k}(\boldsymbol{x})/\sum_{i=1}^{K}\exp v_{i}(\boldsymbol{x})~{},\quad\boldsymbol{v}=\boldsymbol{W}\boldsymbol{z}+\boldsymbol{b}~{},$ (1)
where $\boldsymbol{v}\in\mathbb{R}^{K}$ are the logits output by the final layer with weight matrix $\boldsymbol{W}\in\mathbb{R}^{K\times L}$, bias $\boldsymbol{b}\in\mathbb{R}^{K}$, and pre-logit features $\boldsymbol{z}\in\mathbb{R}^{L}$ as inputs. The classifier is trained by minimising the cross entropy (CE) loss on a finite dataset $\mathcal{D}_{\text{tr}}=\{x^{(n)},y^{(n)}\}_{n=1}^{N}$ drawn from distribution $p_{\text{data}}(\boldsymbol{x},y)$, such that it approximately learns the true conditional $P_{\text{data}}(y|\boldsymbol{x})$
$\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ $=-\frac{1}{N}\sum_{n}\sum_{k}\delta_{y^{(n)}\omega_{k}}\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (2)
$\approx-\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\biggl{[}\sum_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})\log P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})\biggr{]}$ (3)
$=\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\left[\text{KL}\left[\bar{\boldsymbol{\pi}}(\boldsymbol{x})||\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\right]\right]+\text{const.}~{}=\mathcal{L}_{\text{CE}}^{\text{true}}(\boldsymbol{\theta})~{},$ (4)
where $\delta$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence. We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e. $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$. Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$,
$\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})~{}.$ (5)
We also define the probability of the classifier making an error as,
$P_{\text{error}}=1-\bar{\pi}_{\hat{y}}~{},$ (6)
where, in a slight abuse of notation $\bar{\pi}_{\hat{y}}$ is the true probability of the predicted class.
2.1 Label Smoothing (LS)
Figure 2: Left: illustration of how label smoothing (LS) alters a training label. LS reduces data supervision and adds regularisation. Right: illustration of selective classification (SC). Uncertain samples ($U>\tau$) are rejected, with the aim of reducing the number of errors  served.
Label smoothing is most commonly understood (and implemented) as mixing the original one-hot labels (represented by $\delta$ in Eq. 2) with a uniform categorical distribution $\boldsymbol{u}=\boldsymbol{1}/K$ using hyperparameter $\alpha\in[0,1]$ (see Fig. 2). The label smoothing loss is thus
$\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ $=-\frac{1}{N}\sum_{n}\sum_{k}\left[(1-\alpha)\delta_{y^{(n)}\omega_{k}}+\alpha\frac{1}{K}\right]\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (7)
$=(1-\alpha)\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})+\alpha\frac{1}{N}\sum_{n}\left[\text{KL}\left[\boldsymbol{u}||\boldsymbol{\pi}(\boldsymbol{x}^{(n)};\boldsymbol{\theta})\right]\right]+\text{const.}~{}$. (8)
$\approx\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\Bigl{[}\text{KL}\Bigl{[}\underbrace{(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}\,\,+\!\!\!\!\underbrace{\alpha\boldsymbol{u}}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!||\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\Bigr{]}\Bigr{]}+\text{const.}=\mathcal{L}_{\text{LS}}^{\text{true}}(\boldsymbol{\theta};\alpha)~{},$ (9)
where we see that it can also be viewed as reduced CE supervision from the data combined with a regularisation term encouraging the softmax $\boldsymbol{\pi}$ to be uniform and preventing it from overfitting to the training data (Fig. 2). Eq. 9 also shows that LS can be seen as learning to predict a “softened” version of the true conditional $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$.
2.2 Selective Classification (SC)
A simple downstream task for estimates of predictive uncertainty is to reject (or detect) predictions that may incur a high cost [36, 81], using a binary rejection function,
$g(\boldsymbol{x};\tau)=\begin{cases}0\text{ (reject prediction)},&\text{if }U(\boldsymbol{x})>\tau\text{ (uncertain)}\\ 1\text{ (accept prediction)},&\text{if }U(\boldsymbol{x})\leq\tau\text{ (confident)}~{},\end{cases}$ (10)
where $U(\boldsymbol{x})$ is a scalar measure of predictive uncertainty extracted from the prediction model and $\tau$ is the operating threshold. Intuitively, we reject if the model is uncertain. Notably, we are only concerned with the relative rankings of $U$s rather than the absolute values. This differs from model calibration [25, 36], where the value of uncertainty predicted should match the marginal error rate over the data distribution for that value.
In the case where the prediction task is classification and we wish to reject potential misclassifications (), we can use a selective classifier [14, 9] $(f,g)$, which is simply the combination of a classifier $f$ (Eq. 5) and the aforementioned binary rejection function $g$ (Eq. 10). Fig. 2 contains an illustration. Given the 0/1 classification error,
[…]
we define the selective risk [14, 22] as
$\text{Risk}(f,g;\tau)=\frac{\mathbb{E}_{p_{\text{data}}(\boldsymbol{x},y)}[g(\boldsymbol{x};\tau)\mathcal{L}_{\text{SC}}(f(\boldsymbol{x}),y)]}{\mathbb{E}_{p_{\text{data}}(\boldsymbol{x},y)}[g(\boldsymbol{x};\tau)]}~{},$ (12)
[…]
$U(\boldsymbol{x})=-\text{MSP}(\boldsymbol{x})=-\pi_{\text{max}}=-\max_{k}\pi_{k}(\boldsymbol{x};\theta)=-P(\hat{y}|\boldsymbol{x};\boldsymbol{\theta})~{},$ (13)
i.e. the (negative of the) Maximum Softmax Probability (MSP) [30] is the model’s estimate of the probability of its prediction $\hat{y}$. We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature [64, 81, 83, 36, 4, 15, 38, 22, 30].
2.2.1 Overconfidence.
Since we are only concerned about the rank ordering of uncertainties $U$, we loosely refer to overconfidence as when a model’s estimate of uncertainty is (relatively) low when $P_{\text{error}}$ high, as intuitively we would like a model to be uncertain when it is likely to be wrong. Underconfidence is then the inverse. Both will result in worse SC performance as overconfidence will cause errors to be accepted whilst underconfidence will lead to correct predictions being rejected. This is different to the more specific definition used in the problem scenario of model calibration [25].
2.3 Aleatoric and Epistemic Uncertainty
Figure 3: Left: different types of uncertainty for a model trained on ImageNet [63]. Aleatoric uncertainty is high as a different label $y$ could be sampled, while epistemic uncertainty is high as the image is from an unfamiliar domain, here ImageNet-Sketch [76]. Right: how errors arise from different uncertainties. Aleatoric uncertainty leads to errors as there are many probable labels. Epistemic uncertainty leads to errors as the model is unable to predict the true distribution $\bar{\boldsymbol{\pi}}$.
It is useful to separate uncertainty into two different types [39, 19, 35]. Aleatoric (or data) uncertainty is irreducible uncertainty in the data distribution $P_{\text{data}}(y|\boldsymbol{x})$, e.g. when there are multiple valid labels. Epistemic (or knowledge) uncertainty arises from a model’s lack of knowledge about data. A model should have greater epistemic uncertainty on the tails or low-density regions of $p_{\text{data}}(\boldsymbol{x})$, as these regions of the data distribution will be weighted less in the training loss (Eq. 4). There will also be epistemic uncertainty for different images not in training data, or images from a different domain. These different uncertainties are illustrated visually in Fig. 3 (left). While errors caused by aleatoric uncertainty are unavoidable, those due to epistemic uncertainty arise from the model’s inability to accurately predict the true conditional distribution $\bar{\boldsymbol{\pi}}$. These can be avoided with additional supervision. Both types of errors are illustrated in Fig. 3 (right). Epistemic uncertainty also intuitively results in both overconfidence and underconfidence. As the neural network models the data less accurately, its uncertainty estimates also more poorly reflect $P_{\text{error}}$ (both cases in Fig. 3 (right) are overconfident).
2.4 Experimental Details
We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Appendix 0.A. image classification on ImageNet-1k [63] and semantic segmentation (pixel-level classification) on Cityscapes [12]. For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation. For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images. To estimate the risk, we extract the model output prior to the final interpolation and subsample 5000 labelled pixels per image at random. We evaluate on the same pixels between models. All experiments are performed in Pytorch [61] and we will release our code shortly.
2.4.1 Models and Training
The only parameter varied between training runs is the level of label smoothing $\alpha$. Moreover, in order to isolate the effects of label smoothing, we train all our models from scratch using simple recipes. We purposely avoid popular augmentations such as MixUp [87] and CutMix [86] as these directly affect the training labels, which would interfere with our experiments. For image classification, we train ResNet-50 [27] and ViT-S-16 [13] on ImageNet-1k [63] using only random resized cropping and flipping for data augmentation. To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) [16, 8]. For semantic segmentation, we train DeepLabV3+ [7] (ResNet-101 backbone) on Cityscapes [12]. We only use random cropping, flipping and colour jitter for augmentations. Full training details for reproducing our results can be found in the Appendix 0.B.
3 Label Smoothing Degrades Selective Classification Performance
Figure 4: Risk-coverage plots for ViT-S-16-SAM trained on ImageNet-1k and DeepLabV3+ (ResNet-101) trained on Cityscapes, for different levels of LS $\alpha$. Although LS may improve error rate at 100% coverage, increasing $\alpha$ consistently degrades selective classification performance.
Figure 5: A few illustrative examples of overconfident errors performed by our LS-trained ResNet-50. Even though $P_{\text{error}}$ is high in all cases the model has predicted low-ranking (amongst all samples) uncertainty. Note that even though the CE model is wrong as well, it has assigned higher ranking uncertainties, reflecting its superior SC ability shown in Fig. 1. As an aside we note the high aleatoric uncertainty in ImageNet, arising from images containing multiple possible labels.
Figure 6: Visualisation of the overconfidence of DeepLabV3+ (ResNet-101) when trained using LS. We see that compared to the model trained using CE, the LS model predicts much lower uncertainty on incorrect ✗ segmentations. In particular, for the erroneous region on the left where the model has predicted parts of the “sidewalk” as “road”, the LS model is highly overconfident. This could have dire consequences in a safety-critical application such as autonomous driving.
Recently, large-scale empirical evaluations of SC for pre-trained models from open-source repositories [20, 4] have revealed that many strong classifiers surprisingly underperform on SC. In this section, we present LS as one possible cause of these previously unexplained results, and explain how LS may affect a model’s ability to perform SC.
To examine the effects of LS on SC performance, we plot RC curves for different architectures and tasks in Figs. 1 and 4. We use $U=-\text{MSP}$ and vary only the LS level $\alpha$ between training runs. We see that training with LS leads to a noticeable degradation for selective classification, despite improvements in top-1 error rate (risk at 100% coverage). That is to say, LS weakens the MSP score’s ability to differentiate  vs . We provide illustrative examples of this overconfidence on ImageNet in Fig. 5. The degradation in performance is especially evident for the low-risk regime, which is relevant to safety-critical scenarios where tolerances for error are low. For example, if the target is to achieve only 1% error on ImageNet, then none of our LS models are able to achieve a meaningful level of coverage, effectively not being able to serve any predictions.
To further highlight the overconfidence caused by LS, we visualise the uncertainty ranking of incorrect pixels for the segmentation of a Cityscapes scene in Fig. 6. We see that the LS-trained model is extremely confident for an erroneous region where it has predicted the “sidewalk” as “road”. Whilst the CE-trained model has made the same error, it is much more uncertain. This illustrates potential danger in an autonomous driving scenario if the vehicle is making decisions based on uncertainty estimates.
Upon inspection, many of the underperforming models benchmarked in [20, 4] are indeed trained using LS, [^3] We include a more detailed discussion of existing benchmarks in Sec. 0.D.2. further supporting our own results. The models are sourced from repositories such as from torchvision [61] and timm [79], where the original training recipes are mainly optimised for top-1 accuracy. [^4] This blog post is a good example https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/ . We note that in the case of ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g. [72, 73]. Overall, our results highlight that blindly optimising for accuracy may result in negative downstream consequences and that practitioners of selective classification need to be aware of the effects of their training recipes.
3.1 Considering Overconfidence Arising from Epistemic Uncertainty
It might intuitively seem odd that LS leads to such degraded SC performance since the transform applied to the true targets in Eq. 9 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$ does not change the relative ranking of the maximum probability. However, this is only true for low epistemic uncertainty, i.e., when the model is knowledgeable about the data, so $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$. For such samples, the model has learnt to be more uncertain. However, we cannot assume that epistemic uncertainty is always low. Moreover, we know it leads to both errors and overconfidence (Fig. 3 (right)). One clue that points to epistemic uncertainty is the fact that for the LS-trained models, Figs. 1 and 4 show that many samples have $\pi_{\text{max}}>(1-\alpha)+\alpha/K$, and it is for these such samples that SC performance is degraded the most. As the training targets never exceed $(1-\alpha)+\alpha/K$, this inaccuracy in predicting the smoothed targets could be due to epistemic uncertainty.
3.2 Comparing Logit Gradients Between CE and LS
To better understand how LS could lead to the poor SC performance exhibited in the RC curves of Figs. 1 and 4, we consider how LS affects logit-level training gradients. These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training. We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs. 4 and 9), which the empirical loss approximates, for a single sample,
$\frac{\partial\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=-\left[\bar{\pi}_{k}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\bar{\pi}_{k}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (14)
where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the regularisation gradient on the logits,
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (15)
which is the difference between the LS and CE gradients. This intuitively represents how LS regularises training at the logit level in comparison to CE. Crucially, it only depends on the true distribution $\bar{\boldsymbol{\pi}}$. Gradient descent involves updating weights in the opposite direction to the gradient. Hence $\alpha\bar{\pi}_{k}$ penalises $v_{k}$ when the true probability $\bar{\pi}_{k}$ is higher. The second term $-\alpha/K$ uniformly increases the logits. This does not affect the softmax output as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$ [44].
3.3 Label Smoothing Exacerbates Overconfidence and Underconfidence
Figure 7: Illustration of how the regularisation gradient (Eq. 15), which is the difference between LS and CE gradients, affects the logits differently. When $U$ is lower (sharper softmax), the regularisation on the max logit is lower when epistemic uncertainty has caused overconfidence. When $U$ is higher (flatter softmax), the regularisation is higher when epistemic uncertainty has caused underconfidence. Thus, LS exacerbates both of these issues, degrading the rank ordering of $U$ and diminishing the ability of the softmax to differentiate correct  from incorrect  .
We now consider how the regularisation gradient affects the maximum logit,
$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (16)
[…]
To further build an intuition, we consider overconfidence and underconfidence as a result of epistemic uncertainty. As illustrated in Fig. 7, when $U$ is low, and a model is overconfident due to epistemic uncertainty (i.e. $P_{\text{error}}$ is high), the regularisation gradient is much weaker compared to when $P_{\text{error}}$ is low for the same $U$. On the other hand, when $U$ is high and a model is underconfident due to epistemic uncertainty (i.e. $P_{\text{error}}$ is low), the regularisation gradient is much stronger compared to when $P_{\text{error}}$ is high for the same $U$. Thus, both overconfidence and underconfidence are exacerbated by label smoothing, degrading the model’s ability to perform selective classification.
To test our analysis, we artificially inject epistemic uncertainty into our ImageNet evaluation using ImageNet-Sketch [76], a dataset containing sketches of each ImageNet class (see Fig. 3 for an example). We combine the 50,889 ImageNet-Sketch images with the 50,000 ImageNet evaluation images and plot the new risk vs coverage for all 100,889 samples in Fig. 8. We find that even though the regularisation of LS still improves the error rate, the degradation in SC from LS is exacerbated at lower coverages, where the CE-trained model maintains much lower risk. At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch. This demonstrates that LS exacerbates overconfidence arising from epistemic uncertainty, reducing a model’s ability to reject errors.
@10% coverage of ImageNet + Sketch
ImageNet
Sketch
CE
#samples
8994
1094
#errors
93
80
error rate
1.0
7.3
LS $\alpha=0.1$
#samples
8561
1527
#errors
187
211
error rate
2.2
13.8
LS $\alpha=0.2$
#samples
8171
1917
#errors
244
365
error rate
3.0
19.0
LS $\alpha=0.3$
#samples
7915
2173
#errors
297
541
error rate
3.8
24.9
Figure 8: Left: RC curve for ResNet-50 when evaluation is on the combination of ImageNet and ImageNet-sketch (ViT-S-16 can be found in Appendix 0.A). We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when epistemic uncertainty is artificially introduced. Right: statistics @10% coverage of the combined evaluation set. As the level of LS $\alpha$ increases, the number of errors increases, especially the number of errors from ImageNet-Sketch. This shows that LS exacerbates model overconfidence arising from epistemic uncertainty.
4 Logit Normalisation Improves the SC of LS-Trained Models
Ideally, we would like to find a way to recover from the degradation caused by LS. A recent study by Cattelan and Silva [4] has shown that logit normalisation can improve the SC performance of many pre-trained models. The logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit
$U=-v^{\prime}_{\text{max}},\quad\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}}{\left\lVert\boldsymbol{v}\right\rVert_{p}}=\boldsymbol{v}/\bigl{(}\sum_{k}v_{k}^{p}\bigr{)}^{\frac{1}{p}}~{},$ (17)
where $p$ is found via grid search on a validation set. We investigate the efficacy of this approach, applying it to our LS-trained models. Fig. 9 (left) shows the SC performance with and without logit normalisation, and we indeed find that applying logit normalisation greatly improves SC performance for models trained using LS, allowing for improved error rate (@100% coverage) without compromising on SC. This is further visualised for semantic segmentation in Fig. 10, where logit normalisation successfully mitigates the previously observed overconfidence caused by LS. However, we also find that logit normalisation does not notably improve the ability of CE-trained models. This aligns with the results reported [4] where certain models do not benefit from logit normalisation so the authors suggest “falling back” to the MSP.
Figure 9: Left: RC curves showing the effect of logit normalisation. Logit normalisation improves SC performance on LS-trained models, but has little effect on CE-trained models. DeepLabV3+ can be found in Appendix 0.A. Right: distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately. We see that $v_{\text{max}}$ is higher for  for the LS-trained model, whilst the distributions are roughly the same for CE. This matches the max logit regularisation described in Eq. 16. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
Figure 10: Visualisation of the effect of logit normalisation for semantic segmentation (same scene as Fig. 6). Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model.
4.1 Explaining the Effectiveness of Logit Normalisation.
Although Cattelan and Silva [4] validate this approach on a large number of pre-trained models, they provide little explanation as to why it is so effective (or indeed why it isn’t effective sometimes). Let us highlight the interaction between logit normalisation and label smoothing. If we examine Eq. 17, we see that it resembles the softmax. However, in this case, the logits are raised to power $p$ rather than being exponentiated. Crucially, whilst the softmax is invariant to uniform shifts in the logits,
$\pi_{k}(\boldsymbol{v})=\frac{\exp v_{k}}{\sum_{i}\exp v_{i}}=\frac{\exp(v_{k}+\eta)}{\sum_{i}\exp(v_{i}+\eta)}=\pi_{k}(\boldsymbol{v}+\eta\boldsymbol{1}),\quad\forall\eta\in\mathbb{R}~{},$ (18)
this is not the case for $\boldsymbol{v}^{\prime}$. In fact, we have the following inequality for positive [^5] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits. We include additional discussion in Sec. 0.D.1. logits $\boldsymbol{v}$:
Result 1.
For all strictly positive vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$, containing at least two different values, and $p\in[1,+\infty[$, the ratio of the $\infty$-norm and the $p$-norm strictly decreases when summing $v$ with any uniform vector $\eta\boldsymbol{1}$, $\eta$ strictly positive:
$\frac{||\boldsymbol{v}||_{\infty}}{||\boldsymbol{v}||_{p}}>\frac{||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}}{||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}}~{}.$ (19)
The proof can be found in Appendix 0.C. Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of $v^{\prime}_{\text{max}}$. That is to say logit normalisation increases the estimated uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$.
Recall that in Sec. 3.3 we found that the maximum logit $v_{\text{max}}$ is less regularised the more likely a prediction is wrong ✗, i.e. when $P_{\text{error}}$ is high (Eq. 16). We plot $v_{\text{max}}$ given the MSP, separately for correct ✓ and incorrect ✗ predictions in Fig. 9, and see an interesting behaviour emerge. For the model trained with LS, the imbalanced logit regularisation described in Eq. 16 has caused $v_{\text{max}}$ to become higher for misclassifications ✗ than correct predictions ✓, i.e. information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$.
As Result 1 says that for the same softmax output $\boldsymbol{\pi}$, $v^{\prime}_{\text{max}}$ will decrease as $v_{\text{max}}$ increases, logit normalisation will increase uncertainty when $v_{\text{max}}$ is higher, effectively reversing the effect of the imbalanced logit regularisation. This will then improve the rank ordering of the uncertainty score $U$ and thus improve the SC performance of the LS-trained model. Given our analysis and empirical results, we strongly recommend the use of logit normalisation for LS-trained models for performing selective classification.
On the other hand, Fig. 9 (right) also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE-trained model. This explains why, for the models trained using CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.
5 Related Work
5.0.1 Prediction with Rejection.
Selective classification falls into the broader problem setting of prediction with rejection. In the case of SC, misclassifications are to be rejected [14]. The baseline approach is to use the MSP [30, 22] and there have been a number of proposed training [57, 33, 89] and architectural [23, 11] enhancements, however, recently the effectiveness of some of these enhancements has been called into question [15]. In this work, we focus on only the MSP. Another scenario involving prediction with rejection is out-of-distribution (OOD) detection [84, 30], where data from outside of the training distribution are to be rejected. There is a plethora of varied research in this field [82, 68, 49, 88, 50, 34, 77, 29] extending to semantic segmentation [26, 46, 2, 24, 18]. Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data [36, 81, 38, 5]. Notably, Deep Ensembles [43] have arisen as a reliable method to improve performance in all three scenarios [38, 83, 45]. We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
5.0.2 Mixup.
Mixup [87] and its variants [86, 17, 62, 54, 3] are a set of regularisation techniques that involve interpolating between random pairs of samples at training time, modifying both inputs and targets. They are commonly found in many ImageNet training recipes [73, 52, 51] and often used in conjunction with label smoothing [80, 74, 53]. We remark that research into how Mixup and its variants affect SC is a particularly salient avenue of future work as it similarly “softens” the classification training targets.
5.0.3 Label Smoothing.
Beyond prediction with rejection, label smoothing [70] is well investigated. It has been shown to improve model calibration [59, 58, 47, 10] as well as accuracy when training under label noise [56]. Knowledge distillation [31], where soft labels are provided by a teacher network, is also commonly linked and combined with LS [59, 21, 65, 6, 85] due to their similarity. Interestingly, in a similar vein to our work, pre-training using LS has been shown to harm transfer learning [40].
6 Concluding Remarks
In this work, we investigate the effect that label smoothing has on a model’s ability to perform selective classification. Our experiments across a range of architectures and tasks reveal that the use of label smoothing leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves classification performance. By analysing the logit-level gradients of the label smoothing loss, we are able to explain this behaviour – label smoothing increasingly regularises the max logit as the true probability of error decreases, exacerbating overconfidence and underconfidence. We then investigate logit normalisation as a potential method to improve the degraded selective performance of label-smoothing-trained models. We find it to be highly effective, as it, in essence, reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher. We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
6.0.1 Acknowledgments.
This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-[AD011011970R3]). GX’s PhD is funded jointly by Arm and the EPSRC.
Supplementary Material
Appendix 0.A Additional Results
We include additional experimental results that mirror those found in the main paper on ResNet-50, for ViT-S-16 and DeepLabV3+:
  • Fig. 11 shows ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch. The behaviour mirrors that of ResNet-50 in the main paper.
  • Fig. 12 shows the effectiveness of logit normalisation on DeepLabV3+ (ResNet-101) on Cityscapes.
  • Figs. 13 and 14 shows the distribution of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ViT-S-16 and DeepLabV3+. We see that similarly to ResNet-50, for LS, the distribution of $v_{\text{max}}$ is higher for errors ✗. Although it is less obvious, it is clear for both ViT-S-16 and DeepLabV3+ that for higher $\pi_{\text{max}}$ the standard deviations overlap much less than for CE.
@10% coverage of ImageNet + Sketch
ImageNet
Sketch
CE
#samples
9528
560
#errors
88
43
error rate
0.9
7.7
LS $\alpha=0.1$
#samples
9332
756
#errors
130
83
error rate
1.4
11.0
LS $\alpha=0.2$
#samples
9117
971
#errors
177
164
error rate
1.9
16.9
LS $\alpha=0.3$
#samples
8966
1122
#errors
208
195
error rate
2.3
17.4
Figure 11: ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch. The results mirror ResNet-50 from the main paper. LS leads to increased overconfidence when epistemic uncertainty is introduced, as shown through the increase in the failure rate amongst high-confidence samples.
Figure 12: RC curve showing the effect of logit normalisation for DeepLabV3+ (ResNet-101). The behaviour is similar to the results in the main paper. Logit normalisation is effective in improving the performance of the LS-trained model, however, makes little difference to the CE-trained model.
Figure 13: The max logit $v_{\text{max}}$ as a function of the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16. Similarly to the main paper, the distribution is higher for errors  for the LS-trained model. We calculate the mean±$\pm$±std. in a 0.05-wide sliding window.
Figure 14: The same as Fig. 13 but for DeepLabV3+ (ResNet-101)
0.A.1 Small-Scale CIFAR Experiments
We also include experimental results on small-scale $32\times 32$ CIFAR-100 [41]. We train a DenseNet-BC [32] ($k=12,L=100$) to show further generality over model architecture families. Figs. 15, 16 and 17 show results that mirror those found in the main paper for ResNet-50 on ImageNet, although we note that LS does not improve top-1 error rate in this case.
Figure 15: RC curves for DenseNet on CIFAR-100 – LS degrades SC
Figure 16: DenseNet on CIFAR-100 – logit normalisation improves SC for LS but not for CE
Figure 17: DenseNet on CIFAR-100 – LS leads to higher $v_{\text{max}}$ on errors
Appendix 0.B Reproducibility
Alongside this document, we provide a code demo to train two ResNet-20 [27] on CIFAR-10 [41] with cross-entropy and label smoothing and compare the corresponding Risk-Coverage curves. We recall that all the code used in this project will be made available shortly to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix 0.A. All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
0.B.1 Image Classification
DenseNet – CIFAR100.
For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set. We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov [60, 69], a momentum of 0.9, and $1\times 10^{-4}$ weight decay. We divide the learning rate by ten after 150 and 225 epochs. We use standard augmentations. We apply random crop with a four-pixel padding as well as random horizontal flip. We do not perform model selection and keep the last checkpoint.
ViT-16-S – ImageNet.
For our ViT-16-S trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW [55] with the $\beta$s equal to 0.9 and 0.999. We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate. The models are trained with non-adaptive sharpness-aware minimisation (SAM) [16, 8] with $\rho=0.2$. We use a dropout [66] rate of 0.1 but no attention dropout. We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip. For evaluation, we center-crop the images to this resolution. For ImageNet, we do not perform model selection and keep the last checkpoint. However, we randomly extract a validation set of 50000 images from the training set to perform logit normalisation.
ResNet-50 – ImageNet.
Our ResNet-50 is trained on 8 V100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images. After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We use the same transformation of the images as for the ViT and select the last checkpoint for inference. We use the same evaluation set as for the ViT for logit normalisation.
0.B.2 Semantic Segmentation
Deeplabv3+.
We train a Deeplabv3+ on CityScapes [12] with a ResNet-101 [27] backbone pre-trained on ImageNet [63]. We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy [48] with a power of 0.9. The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$. We take a batch size of 12 images and train for 40000 steps. During training, we randomly crop the input images and targets to squares with 768-pixel-long sides. We apply random horizontal flip and colour-jitter with the classical parameters: brightness, contrast, and saturation levels of 0.5. For testing, we use the images at their original resolution and do not perform any test time augmentations. For the RC curves, we randomly sample 5000 predictions per image, extracted prior to the final interpolation to reduce correlations between the predictions. We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
Appendix 0.C Result and Proof
Result.
For all strictly positive vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$ containing at least two different values and $p\in[1,+\infty[$, the ratio of the infinite norm and the p-norm strictly decreases when summing $v$ and any uniform vector $\eta\boldsymbol{1}$, $\eta$ strictly positive:
$\frac{||\boldsymbol{v}||_{\infty}}{||\boldsymbol{v}||_{p}}>\frac{||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}}{||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}}.$ (20)
Proof.
Let there be a real $\eta>0$. Take $p\geq 1$ the dimension of the norm and $\boldsymbol{v}$ a vector of dimension $K\geq 1$ of strictly positive elements $v_{k}$ for $1\leq k\leq K$, such that there exists $1\leq i\leq K$ such that $v_{i}<\max\limits_{k\leq K}v_{k}$. We have that
$1+\frac{\eta}{v_{k}}\geq 1+\frac{\eta}{\max\limits_{k\leq K}v_{k}},$ (21)
and, for at least $i$, we have the same equation, yet with strict inequality. We can adapt Eq. 21 to get
$v_{k}+\eta\geq\frac{\max\limits_{k\leq K}(v_{k}+\eta)}{\max\limits_{k\leq K}v_{k}}v_{k}.$ (22)
And when set to exponent $p\geq 1$, we obtain
$\frac{(v_{k}+\eta)^{p}}{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}\geq\frac{v_{k}^{p}}{\max\limits_{k\leq K}v_{k}^{p}}.$ (23)
As for Eq. 21, please note that using $k=i$, we get the same equation as Eq. 23, although with a strict inequality. We can now sum on the elements of $\boldsymbol{v}$ to get
$\frac{\sum\limits_{k=1}^{K}(v_{k}+\eta)^{p}}{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}>\frac{\sum\limits_{k=1}^{n}v_{k}^{p}}{\max\limits_{k\leq K}v_{k}^{p}}.$ (24)
By taking the inverse, we get
$\frac{\max\limits_{k\leq K}v_{k}^{p}}{\sum\limits_{k=1}^{n}v_{k}^{p}}>\frac{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}{\sum\limits_{k=1}^{K}(v_{k}+\eta)^{p}}.$ (25)
And setting the equation to the exponent $p^{-1}$ and replacing the maxima of the $v_{k}$ and $v_{k}+\eta$ with the infinite norm, $||\boldsymbol{v}||_{\infty}$ and $||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}$ respectively, we obtain the result. ∎
Appendix 0.D Discussions
0.D.1 On the Assumption that Logits are Positive
Figure 18: Mean±$\pm$±std after logits have been sorted highest to lowest for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. We see that $v^{5}$ and $\exp v$ are much larger for the top $<10$ logits. Thus these logits dominate $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$.
Figure 19: $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ within bins of similar MSP for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. Generally, $v^{\prime}_{\text{max}}$ increases as $v_{\text{max}}$ decreases, showing empirically how logit normalisation increases uncertainty for higher $v_{\text{max}}$. We calculate the mean±$\pm$±std. of $v^{\prime}_{\text{max}}$ in a 0.2-wide sliding window for samples with $v_{\text{max}}$ within mean±$\pm$2std of $v_{\text{max}}$ within the bin (to remove noisy averages).
Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true. However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits. This is intuitive as exponentiating or raising to power $p$ ($p\geq 4$ in our experiments) will amplify the larger logits. This is shown in Fig. 18, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive. As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to penalise higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
We can consider the scenario where we add $\eta$ to only the top-$k$ logits. This more aptly describes the empirical logit behaviour compared to adding $\eta$ to all logits, as the values of the lower ranking logits vary much less than the higher ranking ones (Fig. 18). Here we would expect $\pi_{\text{max}}(\boldsymbol{v})$ to increase very slightly, but would expect $v^{\prime}_{\text{max}}(\boldsymbol{v})$ to decrease as the numerator of Eq. 24 would be dominated by the top-$k$ largest logits.
Fig. 19 shows how empirically logit normalisation indeed increases uncertainty for higher $v_{\text{max}}$. We plot the mean±$\pm$±std of $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ for samples in different MSP bins. We see clearly that in almost all cases, for samples with similar $\pi_{\text{max}}$, the normalised max logit $v^{\prime}_{\text{max}}$ increases as the original max logit $v_{\text{max}}$ decreases. As such, logit normalisation is able to improve the SC performance of LS-trained models.
0.D.2 Existing Benchmarks and Training Recipes
Although we do not exhaustively search all training recipes for all models benchmarked in [20, 4], we do provide a number of examples of evaluated models trained with label smoothing. We also provide links to publicly available training repositories, as not all papers mention label smoothing even when it is used in training. Upon inspection of [20, 4], these models do in fact seem to underperform at selective classification (and Cattelan and Silva [4] report that their AURCs benefit from logit normalisation).
Galil et al. [20] state that some of their best performing ViT models [13, 67, 8] are trained with label smoothing (their Tab.1). However, after inspecting both the original papers and open-source repositories [^6] https://github.com/google-research/vision_transformer of the aforementioned work we were unable to find any confirmation of the use of label smoothing.

### v2

@@
-Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It
+Towards Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It
 Abstract
-Label smoothing (LS) is a popular regularisation method for training deep neural network classifiers due to its effectiveness in improving test accuracy and its simplicity in implementation.
+Label smoothing (LS) is a popular regularisation method for training neural networks as it is effective in improving test accuracy and is simple to implement.
 “Hard” one-hot labels are “smoothed” by uniformly distributing probability mass to other classes, reducing overfitting.
-In this work, we reveal that LS negatively affects selective classification (SC) – where the aim is to reject misclassifications using a model’s predictive uncertainty.
-We first demonstrate empirically across a range of tasks and architectures that LS leads to a consistent degradation in SC.
-We then explain this by analysing logit-level gradients, showing that LS exacerbates overconfidence and underconfidence by regularising the max logit more when the probability of error is low, and less when the probability of error is high.
+Prior work has suggested that in some cases LS can degrade selective classification (SC) – where the aim is to reject misclassifications using a model’s uncertainty.
+In this work, we first demonstrate empirically across an extended range of large-scale tasks and architectures that LS consistently degrades SC.
+We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by regularising the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong.
 This elucidates previously reported experimental results where strong classifiers underperform in SC.
-We then demonstrate the empirical effectiveness of logit normalisation for recovering lost SC performance caused by LS.
-Furthermore, based on our gradient analysis, we explain why such normalisation is effective.
-We will release our code shortly.
-Keywords:
-Uncertainty Estimation Selective Classification Label Smoothing
+We then demonstrate the empirical effectiveness of post-hoc logit normalisation for recovering lost SC performance caused by LS.
+Furthermore, linking back to our gradient analysis, we again provide an explanation for why such normalisation is effective.
 1 Introduction
-Label smoothing (LS) [70] is a common regularisation technique used to improve classification accuracy in supervised deep learning.
-When training with cross-entropy minimisation, one-hot labels are linearly combined with a uniform distribution over classes, redistributing the probability mass and “smoothing” the “hard” targets.
-This discourages overfitting on the training labels.
-Due to its simplicity and empirical effectiveness, label smoothing features in many recent training recipes [75, 71, 28, 74, 52, 51, 72, 73, 53], being particularly popular on the ImageNet-1k [63] image classification benchmark.
-Within the domain of uncertainty estimation, label smoothing is well explored in the context of model calibration [59, 10, 58, 47], where the aim is to align a model’s output probabilities with its empirical accuracy.
-The pairing is intuitive as label smoothing encourages models to be less confident, and models are typically more confident than they are accurate [25].
-On the other hand, there is a distinct lack of research investigating the effects of label smoothing in the context of Selective Classification.
-Selective Classification (SC) [30, 22, 81, 36] is a problem setting where, in addition to the primary classification task, a binary rejection decision is made based on the predictive uncertainty estimated by the model, i.e.
-reject if uncertain.
-The aim is to reduce the number of errors served by the classifier by pre-emptively rejecting potential misclassifications.
-It is well motivated by applications where safety and reliability are important due to the high cost of failure, such as autonomous driving [37] and medical diagnosis [1, 42].
-[…]
-Recent large-scale empirical evaluations of pre-trained models [20, 4] have shown that many strong classifiers have surprisingly poor SC ability.
-That is to say, in spite of having very low error rates, they are poor at rejecting their own misclassifications.
-Unfortunately, by evaluating pre-trained models from open-source repositories, the authors find it difficult to disentangle any potential causes of this behaviour, as model training recipes vary wildly from checkpoint to checkpoint.
-In this work, we reveal that label smoothing is one such potential cause and present the following key contributions:
+Label smoothing (LS) (Szegedy et al., 2016) is a common regularisation technique used to improve classification accuracy in deep learning.
+The one-hot labels used for cross entropy (CE) are linearly combined with a uniform distribution over classes, redistributing the probability mass and “smoothing” the “hard” targets, discouraging overfitting.
+Due to its simplicity and empirical effectiveness, label smoothing features in many recent training recipes (Vaswani et al., 2017; Tan et al., 2019; He et al., 2019; Touvron et al., 2021; Liu et al., 2021; 2022b; 2022c; Tan & Le, 2019; 2021), being particularly popular on the ImageNet-1k (Russakovsky et al., 2015) image classification benchmark.
+Within the domain of uncertainty estimation, LS is well explored in the context of model calibration (Müller et al., 2019; Chun et al., 2020; Mukhoti et al., 2020; Liu et al., 2022a), where the aim is to align a model’s output probabilities with its empirical accuracy.
+The pairing is intuitive as LS encourages models to output lower probabilities, and models are typically more confident than they are accurate (Guo et al., 2017).
+On the other hand, there is very little research investigating the effects of LS in the context of Selective Classification.
+Selective Classification (SC) (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017; Xia & Bouganis, 2022b; Jaeger et al., 2023) is a problem setting where, in addition to the primary classification task, a binary rejection decision is made based on the uncertainty estimated by the model, i.e.
+reject/abstain if uncertain.
+The aim is to reduce the number of failures served by the classifier by pre-emptively rejecting potential misclassifications.
+It is well motivated by applications where safety and reliability are important due to the high cost of failure.
+For example, when uncertain, an autonomous driving system may require driver assistance  (Kendall & Gal, 2017), a medical diagnosis system may defer to a doctor (Beam & Kompa, 2021; Kurz et al., 2022), or a visual aid system for the visually impaired may abstain from answering a query (Whitehead et al., 2022).
+ResNet-50 ImageNet MSP CE LS $\alpha=0.1$ LS $\alpha=0.2$ LS $\alpha=0.3$ LS $\alpha=0.2$ – logit norm
+%coverage@risk=1$\uparrow$ 17.69 0.2 0.53 0.49 21.9
+Figure 1: Top: LS causes overconfidence for semantic segmentation.
+The LS-trained model predicts much lower (ranked) uncertainty on incorrect ✗ segmentations than CE.
+In particular, for the erroneous region on the left where the model has predicted parts of the “sidewalk” as “road”, the LS model is highly overconfident.
+This could have dire consequences in a safety-critical application such as autonomous driving.
+Bottom: LS leads to close to 0% of samples being accepted (coverage) when a strict tolerance of 1% error on accepted samples (risk) is imposed on ImageNet.
+Deployment-time logit normalisation effectively negates the degradation caused by LS.
+Recently Zhu et al.
+(2022; 2024) empirically observe that LS can lead to worse SC for convolutional neural networks (CNNs) performing image classification, yet it is not clear why this occurs.
+Concurrently, large-scale empirical evaluations of open-source pre-trained models have shown that many strong classifiers have surprisingly poor SC ability (Galil et al., 2023; Cattelan & Silva, 2024).
+Cattelan & Silva (2024) additionally present deployment-time logit normalisation as a sometimes-effective approach to improving SC, but the reason for its effectiveness remains unclear.
+In this work, we aim to empirically validate and analytically demystify the effect of LS on SC, tying together and filling in the gaps in knowledge of previous work with the following key contributions:
   1. 1.
-    We show empirically, across a range of network architectures and vision tasks, that training with LS consistently leads to degraded SC performance, even if it may improve accuracy (see Fig.
+    We show empirically, across a range of large-scale architectures (CNN,ViT) and tasks (image classification, semantic segmentation), that training with LS consistently leads to degraded SC performance (see Fig.
+1), even if it may improve accuracy.
+Moreover, we show that the degradation worsens with stronger LS.
+As LS can be found in the training recipes of many of the models evaluated in (Galil et al., 2023; Cattelan & Silva, 2024), this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
+  2. 2.
+    We address a gap in the understanding of LS by providing an explanation of this behaviour through analysing the logit-level gradients of the LS loss.
+We show that the amount LS regularises/suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing(decreasing) the more likely a prediction is correct(wrong).
+This leads to relatively higher uncertainty on correct predictions and lower uncertainty on misclassifications, degrading the ranking of uncertainties and hurting SC compared to vanilla CE.
+  3. 3.
+    We show that post-hoc logit normalisation (Cattelan & Silva, 2024) is effective in negating the degradation from LS (Fig.
 1).
-As LS can be found in the training recipes of many of the models evaluated in [20, 4], this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
-  2. 2.
-    We are able to explain the above behaviour by analysing the logit-level gradients of the LS loss.
-We show that the amount LS regularises the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing as $P_{\text{error}}$ decreases.
-This exacerbates overconfidence and underconfidence, degrading SC.
-  3. 3.
-    Finally, we demonstrate that logit normalisation proposed by Cattelan and Silva [4] is highly empirically effective in combating specifically the degradation caused by LS.
-Moreover, we are able to explain the effectiveness of logit normalisation through the lens of our gradient-based analysis, showing that normalisation compensates for the imbalanced regularisation by increasing uncertainty as the max logit increases.
-[^1] Although they demonstrate its empirical effectiveness, Cattelan and Silva [4] do not make a connection to label smoothing, and provide limited insight as to why their approaches are effective.
+Moreover, we elucidate this effectiveness through the lens of our gradient-based analysis.
+Linking back to the imbalanced logit regularisation of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
 2 Preliminaries
@@
 Typically the network has a categorical softmax output $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\in[0,1]^{K}$,
-$P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})=\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=\exp v_{k}(\boldsymbol{x})/\sum_{i=1}^{K}\exp v_{i}(\boldsymbol{x})~{},\quad\boldsymbol{v}=\boldsymbol{W}\boldsymbol{z}+\boldsymbol{b}~{},$ (1)
+$P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})=\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=\exp v_{k}(\boldsymbol{x})/\sum\nolimits_{i=1}^{K}\exp v_{i}(\boldsymbol{x})~{},\quad\boldsymbol{v}=\boldsymbol{W}\boldsymbol{z}+\boldsymbol{b}~{},$ (1)
 where $\boldsymbol{v}\in\mathbb{R}^{K}$ are the logits output by the final layer with weight matrix $\boldsymbol{W}\in\mathbb{R}^{K\times L}$, bias $\boldsymbol{b}\in\mathbb{R}^{K}$, and pre-logit features $\boldsymbol{z}\in\mathbb{R}^{L}$ as inputs.
-The classifier is trained by minimising the cross entropy (CE) loss on a finite dataset $\mathcal{D}_{\text{tr}}=\{x^{(n)},y^{(n)}\}_{n=1}^{N}$ drawn from distribution $p_{\text{data}}(\boldsymbol{x},y)$, such that it approximately learns the true conditional $P_{\text{data}}(y|\boldsymbol{x})$
-$\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ $=-\frac{1}{N}\sum_{n}\sum_{k}\delta_{y^{(n)}\omega_{k}}\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (2)
-$\approx-\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\biggl{[}\sum_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})\log P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})\biggr{]}$ (3)
+The neural network is trained by minimising the cross entropy (CE) loss on a finite dataset $\mathcal{D}_{\text{tr}}=\{x^{(n)},y^{(n)}\}_{n=1}^{N}$ drawn from distribution $p_{\text{data}}(\boldsymbol{x},y)$, such that it approximately learns the true conditional $P_{\text{data}}(y|\boldsymbol{x})$,
+$\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ $=-\frac{1}{N}\sum\nolimits_{n}\sum\nolimits_{k}\delta_{y^{(n)}\omega_{k}}\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (2)
+$\approx-\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\biggl{[}\sum\nolimits_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})\log P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})\biggr{]}$ (3)
 $=\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\left[\text{KL}\left[\bar{\boldsymbol{\pi}}(\boldsymbol{x})||\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\right]\right]+\text{const.}~{}=\mathcal{L}_{\text{CE}}^{\text{true}}(\boldsymbol{\theta})~{},$ (4)
@@
 $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$.
-Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$,
-$\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})~{}.$ (5)
-We also define the probability of the classifier making an error as,
-$P_{\text{error}}=1-\bar{\pi}_{\hat{y}}~{},$ (6)
-where, in a slight abuse of notation $\bar{\pi}_{\hat{y}}$ is the true probability of the predicted class.
+Predictions $\hat{y}$ are then made on new unlabelled input data $\boldsymbol{x}^{*}$ using classifier function $f$ during deployment,
+$\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}\nolimits_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})~{}.$ (5)
+We also define the probability of the classifier making an error on a given sample as $P_{\text{error}}=1-\bar{\pi}_{\hat{y}}~{}$, where in a slight abuse of notation $\bar{\pi}_{\hat{y}}$ is the true probability of the predicted class $P_{\text{data}}(\hat{y}|\boldsymbol{x})$.
 2.1 Label smoothing (LS)
 Figure 2: Left: illustration of how label smoothing (LS) alters a training label.
-LS reduces data supervision and adds regularisation.
+LS reduces data supervision and adds regularisation, potentially improving generalisation by reducing overfitting.
 Right: illustration of selective classification (SC).
-Uncertain samples ($U>\tau$) are rejected, with the aim of reducing the number of errors  served.
-Label smoothing is most commonly understood (and implemented) as mixing the original one-hot labels (represented by $\delta$ in Eq.
-2) with a uniform categorical distribution $\boldsymbol{u}=\boldsymbol{1}/K$ using hyperparameter $\alpha\in[0,1]$ (see Fig.
+Uncertain samples ($U>\tau$) are rejected/detected, to reduce the number of errors  served by the system.
+Rejected samples can be discarded or processed separately (e.g.
+deferred to a human expert).
+We wish to better separate/rank  vs  via $U$.
+Label smoothing involves mixing the original one-hot labels ($\delta$ in Eq.
+2) with a uniform categorical distribution $\boldsymbol{u}=1/K\cdot\boldsymbol{1}$ using hyperparameter $\alpha\in[0,1]$ (see Fig.
 2).
-The label smoothing loss is thus
-$\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ $=-\frac{1}{N}\sum_{n}\sum_{k}\left[(1-\alpha)\delta_{y^{(n)}\omega_{k}}+\alpha\frac{1}{K}\right]\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (7)
-$=(1-\alpha)\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})+\alpha\frac{1}{N}\sum_{n}\left[\text{KL}\left[\boldsymbol{u}||\boldsymbol{\pi}(\boldsymbol{x}^{(n)};\boldsymbol{\theta})\right]\right]+\text{const.}~{}$.
+The LS loss is thus
+$\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ $=-\frac{1}{N}\sum\nolimits_{n}\sum\nolimits_{k}\left[(1-\alpha)\delta_{y^{(n)}\omega_{k}}+\alpha\frac{1}{K}\right]\log P(\omega_{k}|\boldsymbol{x}^{(n)};\boldsymbol{\theta})$ (6)
+$=(1-\alpha)\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})+\alpha\frac{1}{N}\sum\nolimits_{n}\left[\text{KL}\left[\boldsymbol{u}||\boldsymbol{\pi}(\boldsymbol{x}^{(n)};\boldsymbol{\theta})\right]\right]+\text{const.}~{}$.
 (7)
@@
 Eq.
-9 also shows that LS can be seen as learning to predict a “softened” version of the true conditional $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$.
+8 also shows that LS can be seen as learning to predict a “softened” version of the true conditional $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$, encouraging a model to be less confident on all samples.
 2.2 Selective classification (SC)
-A simple downstream task for estimates of predictive uncertainty is to reject (or detect) predictions that may incur a high cost [36, 81], using a binary rejection function,
+A simple downstream task for estimates of predictive uncertainty is to reject (or detect) predictions that may incur a high cost (Xia & Bouganis, 2022b; Jaeger et al., 2023), using a binary rejection function,
 $g(\boldsymbol{x};\tau)=\begin{cases}0\text{ (reject prediction)},&\text{if }U(\boldsymbol{x})>\tau\text{ (uncertain)}\\ 1\text{ (accept prediction)},&\text{if }U(\boldsymbol{x})\leq\tau\text{ (confident)}~{},\end{cases}$ (9)
-where $U(\boldsymbol{x})$ is a scalar measure of predictive uncertainty extracted from the prediction model and $\tau$ is the operating threshold.
-Intuitively, we reject if the model is uncertain.
-Notably, we are only concerned with the relative rankings of $U$s rather than the absolute values.
-This differs from model calibration [25, 36], where the value of uncertainty predicted should match the marginal error rate over the data distribution for that value.
-In the case where the prediction task is classification and we wish to reject potential misclassifications (), we can use a selective classifier [14, 9] $(f,g)$, which is simply the combination of a classifier $f$ (Eq.
+where $U(\boldsymbol{x})$ is a scalar measure of predictive uncertainty extracted from the prediction model and $\tau$ is a user-set operating threshold.
+Intuitively, we reject if the model is uncertain, preventing the system from serving failures, which may then be processed separately.
+We are only concerned with the relative rankings of $U$s rather than absolute values - we want to be more uncertain on failures.
+In the case where the prediction task is classification and we wish to reject potential misclassifications (✗), we can use a selective classifier (Chow, 1970; El-Yaniv & Wiener, 2010) $(f,g)$, which is simply the combination of a classifier $f$ (Eq.
 5) and the aforementioned binary rejection function $g$ (Eq.
-10).
+9).
 Fig.
 2 contains an illustration.
-Given the 0/1 classification error,
+To evaluate a selective classifier, we use the 0/1 classification error,
 […]
-we define the selective risk [14, 22] as
+to define the selective risk (El-Yaniv & Wiener, 2010; Geifman & El-Yaniv, 2017) as
 $\text{Risk}(f,g;\tau)=\frac{\mathbb{E}_{p_{\text{data}}(\boldsymbol{x},y)}[g(\boldsymbol{x};\tau)\mathcal{L}_{\text{SC}}(f(\boldsymbol{x}),y)]}{\mathbb{E}_{p_{\text{data}}(\boldsymbol{x},y)}[g(\boldsymbol{x};\tau)]}~{},$ (11)
 […]
-$U(\boldsymbol{x})=-\text{MSP}(\boldsymbol{x})=-\pi_{\text{max}}=-\max_{k}\pi_{k}(\boldsymbol{x};\theta)=-P(\hat{y}|\boldsymbol{x};\boldsymbol{\theta})~{},$ (13)
+$U(\boldsymbol{x})=-\text{MSP}(\boldsymbol{x})=-\pi_{\text{max}}=-\max\nolimits_{k}\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})=-P(\hat{y}|\boldsymbol{x};\boldsymbol{\theta})~{},$ (12)
 i.e.
-the (negative of the) Maximum Softmax Probability (MSP) [30] is the model’s estimate of the probability of its prediction $\hat{y}$.
-We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature [64, 81, 83, 36, 4, 15, 38, 22, 30].
-2.2.1 Overconfidence.
-Since we are only concerned about the rank ordering of uncertainties $U$, we loosely refer to overconfidence as when a model’s estimate of uncertainty is (relatively) low when $P_{\text{error}}$ high, as intuitively we would like a model to be uncertain when it is likely to be wrong.
+the (negative of the) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017) is the model’s estimate of the probability of its prediction $\hat{y}$.
+We focus on this uncertainty measure $U$ for SC as it is a popular default choice and has been shown to consistently and reliably perform well in the literature (Xia & Bouganis, 2022b; 2023; Jaeger et al., 2023; Feng et al., 2023) compared to alternatives.
+MSP is a natural choice as $P(y|\boldsymbol{x};\boldsymbol{\theta})$ models $P_{\text{data}}(y|\boldsymbol{x})$, and $U=-\max_{k}P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ (or any scalar monotonic to it) provides the optimal risk if the true distribution $P_{\text{data}}(y|\boldsymbol{x})$ is known (Chow, 1970).
+We omit evaluations on other softmax-based $U$ from the main paper as we find them to behave similarly to MSP.
+In line with previous work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Zhu et al., 2024), we find that OOD detection (Yang et al., 2021) scores perform poorly at SC, and also omit them (see Sec.
+E.1 for discussion and additional results).
+Over/Underconfidence.
+Since we are only concerned about the relative ranking of $U$, we loosely refer to overconfidence as when a model’s estimate of uncertainty for a given sample is (relatively) low when $P_{\text{error}}$ high, as we want a model to be uncertain when it is likely to be wrong.
 Underconfidence is then the inverse.
-Both will result in worse SC performance as overconfidence will cause errors to be accepted whilst underconfidence will lead to correct predictions being rejected.
-This is different to the more specific definition used in the problem scenario of model calibration [25].
-2.3 Aleatoric and Epistemic Uncertainty
-Figure 3: Left: different types of uncertainty for a model trained on ImageNet [63].
-Aleatoric uncertainty is high as a different label $y$ could be sampled, while epistemic uncertainty is high as the image is from an unfamiliar domain, here ImageNet-Sketch [76].
-Right: how errors arise from different uncertainties.
-Aleatoric uncertainty leads to errors as there are many probable labels.
-Epistemic uncertainty leads to errors as the model is unable to predict the true distribution $\bar{\boldsymbol{\pi}}$.
-It is useful to separate uncertainty into two different types [39, 19, 35].
-Aleatoric (or data) uncertainty is irreducible uncertainty in the data distribution $P_{\text{data}}(y|\boldsymbol{x})$, e.g.
-when there are multiple valid labels.
-Epistemic (or knowledge) uncertainty arises from a model’s lack of knowledge about data.
-A model should have greater epistemic uncertainty on the tails or low-density regions of $p_{\text{data}}(\boldsymbol{x})$, as these regions of the data distribution will be weighted less in the training loss (Eq.
+Both will result in worse SC: overconfidence leads to errors being accepted, underconfidence to correct predictions being rejected.
+Intuitively, over/underconfidence may arise from poorness of fit, where $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})$ fails to accurately model the true conditional distribution $\bar{\boldsymbol{\pi}}(\boldsymbol{x})$ (Fig.
 4).
-There will also be epistemic uncertainty for different images not in training data, or images from a different domain.
-These different uncertainties are illustrated visually in Fig.
-3 (left).
-While errors caused by aleatoric uncertainty are unavoidable, those due to epistemic uncertainty arise from the model’s inability to accurately predict the true conditional distribution $\bar{\boldsymbol{\pi}}$.
-These can be avoided with additional supervision.
-Both types of errors are illustrated in Fig.
-3 (right).
-Epistemic uncertainty also intuitively results in both overconfidence and underconfidence.
-As the neural network models the data less accurately, its uncertainty estimates also more poorly reflect $P_{\text{error}}$ (both cases in Fig.
-3 (right) are overconfident).
-2.4 Experimental Details
-We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Appendix 0.A.
-image classification on ImageNet-1k [63] and semantic segmentation (pixel-level classification) on Cityscapes [12].
+This is different to the definition used in model calibration (Guo et al., 2017), which, notably, is concerned with marginal properties (averaged over the input distribution) rather than per-sample properties, as well as absolute probabilities rather than relative rankings.
+It is possible to be highly over/underconfident in calibration but optimal for SC and vice versa (Zhu et al., 2024), e.g.
+applying the monotonic transformation $U=-\max_{k}[P_{\text{data}}(\omega_{k}|\boldsymbol{x})]^{1000}$.
+We note that model calibration is an independent task to SC (Jaeger et al., 2023), is well explored in the context of LS (Müller et al., 2019), and is not the focus of this work.
+LS may improve calibration by increasing the uncertainty of both correct ✓ and incorrect ✗ predictions, however, this will not necessarily help SC (Zhu et al., 2024).
+3 The effect of label smoothing on selective classification
+Zhu et al.
+(2022; 2024) observe empirically (as part of a broader investigation) that for a single value of $\alpha$ LS degrades SC for CNN-based image classification.
+In this section we aim to validate this behaviour on a wider range of large-scale tasks and architectures.
+We find that LS leads to consistent degradation in SC, with stronger LS leading to greater degradation.
+We suggest that it may partially explain the findings of recent large-scale empirical evaluations of open-source pre-trained models (Galil et al., 2023; Cattelan & Silva, 2024) where strong classifiers surprisingly underperform on SC.
+Experimental details.
+We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Sec.
+A.3.
+image classification on ImageNet-1k (Russakovsky et al., 2015) and semantic segmentation (pixel-level classification) on Cityscapes (Cordts et al., 2016).
 For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation.
@@
 We evaluate on the same pixels between models.
-All experiments are performed in Pytorch [61] and we will release our code shortly.
-2.4.1 Models and Training
-The only parameter varied between training runs is the level of label smoothing $\alpha$.
-Moreover, in order to isolate the effects of label smoothing, we train all our models from scratch using simple recipes.
-We purposely avoid popular augmentations such as MixUp [87] and CutMix [86] as these directly affect the training labels, which would interfere with our experiments.
-For image classification, we train ResNet-50 [27] and ViT-S-16 [13] on ImageNet-1k [63] using only random resized cropping and flipping for data augmentation.
-To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) [16, 8].
-For semantic segmentation, we train DeepLabV3+ [7] (ResNet-101 backbone) on Cityscapes [12].
-We only use random cropping, flipping and colour jitter for augmentations.
-Full training details for reproducing our results can be found in the Appendix 0.B.
-3 Label Smoothing Degrades Selective Classification Performance
-Figure 4: Risk-coverage plots for ViT-S-16-SAM trained on ImageNet-1k and DeepLabV3+ (ResNet-101) trained on Cityscapes, for different levels of LS $\alpha$.
-Although LS may improve error rate at 100% coverage, increasing $\alpha$ consistently degrades selective classification performance.
-Figure 5: A few illustrative examples of overconfident errors performed by our LS-trained ResNet-50.
-Even though $P_{\text{error}}$ is high in all cases the model has predicted low-ranking (amongst all samples) uncertainty.
-Note that even though the CE model is wrong as well, it has assigned higher ranking uncertainties, reflecting its superior SC ability shown in Fig.
+The only parameter varied between training runs is the level of LS $\alpha$, and to isolate the effects of LS, we train all models from scratch using simple recipes.
+We purposely avoid augmentations such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019) as these directly affect the training labels, which would interfere with our experiments.
+For ImageNet classification, we train ResNet-50 (He et al., 2016) and ViT-S-16 (Dosovitskiy et al., 2021) using only random resized cropping and flipping for data augmentation.
+To achieve decent accuracy for ViT training from scratch without advanced augmentations, we use sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022).
+For semantic segmentation on Cityscapes, we train DeepLabV3+ (Chen et al., 2018) (ResNet-101 backbone) using only random cropping, flipping and colour jitter for augmentations.
+Full training details for reproducibility can be found in the Appendix B.
+We will release our code in the future.
+3.1 Label smoothing degrades selective classification
+Figure 3: Risk-coverage plots for different levels of LS $\alpha$ for different models and tasks (ImageNet classification and Cityscapes semantic segmentation).
+Although LS may improve error rate at 100% coverage, increasing $\alpha$ increasingly and consistently degrades SC performance.
+To examine the effects of LS on SC, we plot RC curves in Fig.
+3.
+We use $U=-\text{MSP}$ and vary only the LS level $\alpha$ between training runs.
+We see that training with LS leads to a noticeable degradation for selective classification, with higher $\alpha$ worsening the effect.
+We provide illustrative examples of LS overconfidence on ImageNet in Fig.
+9.
+Although LS leads to slightly better risk at higher coverage, it quickly becomes worse than CE as coverage is reduced.
+The degradation is especially evident for the low-risk regime, which is relevant to safety-critical scenarios where tolerances for error are strict.
+For example, if the target is to achieve only 1% error/risk on ImageNet, then all of our LS models have close to zero coverage, rendering them effectively useless (Figs.
+1 and 3).
+To further highlight the overconfidence caused by LS, we visualise the uncertainty ranking of incorrect pixels for the segmentation of a Cityscapes scene in Fig.
 1.
-As an aside we note the high aleatoric uncertainty in ImageNet, arising from images containing multiple possible labels.
-Figure 6: Visualisation of the overconfidence of DeepLabV3+ (ResNet-101) when trained using LS.
-We see that compared to the model trained using CE, the LS model predicts much lower uncertainty on incorrect ✗ segmentations.
-In particular, for the erroneous region on the left where the model has predicted parts of the “sidewalk” as “road”, the LS model is highly overconfident.
-This could have dire consequences in a safety-critical application such as autonomous driving.
-Recently, large-scale empirical evaluations of SC for pre-trained models from open-source repositories [20, 4] have revealed that many strong classifiers surprisingly underperform on SC.
-In this section, we present LS as one possible cause of these previously unexplained results, and explain how LS may affect a model’s ability to perform SC.
-To examine the effects of LS on SC performance, we plot RC curves for different architectures and tasks in Figs.
-1 and 4.
-We use $U=-\text{MSP}$ and vary only the LS level $\alpha$ between training runs.
-We see that training with LS leads to a noticeable degradation for selective classification, despite improvements in top-1 error rate (risk at 100% coverage).
-That is to say, LS weakens the MSP score’s ability to differentiate  vs .
-We provide illustrative examples of this overconfidence on ImageNet in Fig.
-5.
-The degradation in performance is especially evident for the low-risk regime, which is relevant to safety-critical scenarios where tolerances for error are low.
-For example, if the target is to achieve only 1% error on ImageNet, then none of our LS models are able to achieve a meaningful level of coverage, effectively not being able to serve any predictions.
-To further highlight the overconfidence caused by LS, we visualise the uncertainty ranking of incorrect pixels for the segmentation of a Cityscapes scene in Fig.
-6.
-We see that the LS-trained model is extremely confident for an erroneous region where it has predicted the “sidewalk” as “road”.
-Whilst the CE-trained model has made the same error, it is much more uncertain.
+The LS-trained model is extremely confident for an erroneous region where it has predicted the “sidewalk” as “road”.
+Whilst the CE model has made the same error, it is much more uncertain.
 This illustrates potential danger in an autonomous driving scenario if the vehicle is making decisions based on uncertainty estimates.
-Upon inspection, many of the underperforming models benchmarked in [20, 4] are indeed trained using LS, [^3] We include a more detailed discussion of existing benchmarks in Sec.
-0.D.2.
-further supporting our own results.
-The models are sourced from repositories such as from torchvision [61] and timm [79], where the original training recipes are mainly optimised for top-1 accuracy.
-[^4] This blog post is a good example https://pytorch.org/blog/how-to-train-state-of-the-art-models-using-torchvision-latest-primitives/ .
-We note that in the case of ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g.
-[72, 73].
-Overall, our results highlight that blindly optimising for accuracy may result in negative downstream consequences and that practitioners of selective classification need to be aware of the effects of their training recipes.
-3.1 Considering Overconfidence Arising from Epistemic Uncertainty
-It might intuitively seem odd that LS leads to such degraded SC performance since the transform applied to the true targets in Eq.
-9 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$ does not change the relative ranking of the maximum probability.
-However, this is only true for low epistemic uncertainty, i.e., when the model is knowledgeable about the data, so $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\boldsymbol{u}$.
+Upon inspection, many of the SC-underperforming models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024) are trained using LS, aligning with our results.
+The models are sourced from repositories such as torchvision (Paszke et al., 2019) and timm (Wightman, 2019), where the training recipes are optimised for top-1 accuracy.
+We note that for ImageNet, LS is such a common technique that it is often used by default, and not even mentioned in papers, e.g.
+(Tan & Le, 2019; 2021).
+Overall, our results highlight that only optimising for accuracy may result in negative downstream consequences and that practitioners of SC need to be aware of the effects of their training recipes.
+We include additional discussion of existing benchmarks and model checkpoints in Sec.
+E.4.
+4 Towards explaining the negative effect of LS on SC
+In the following section, we attempt to delve deeper into why our empirical results (and those in (Zhu et al., 2024)) occur, shedding light on a previously unexplained phenomenon.
+It might intuitively seem odd that LS degrades SC performance as the transform applied to the true targets in Eq.
+8 $(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$ reduces the max probability for all samples but does not change the relative ranking.
+However, this only carries over to the model when it is knowledgeable about the data, so it is well fit $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\approx(1-\alpha)\bar{\boldsymbol{\pi}}(\boldsymbol{x})+\alpha\boldsymbol{u}$.
 For such samples, the model has learnt to be more uncertain.
-However, we cannot assume that epistemic uncertainty is always low.
-Moreover, we know it leads to both errors and overconfidence (Fig.
-3 (right)).
-One clue that points to epistemic uncertainty is the fact that for the LS-trained models, Figs.
-1 and 4 show that many samples have $\pi_{\text{max}}>(1-\alpha)+\alpha/K$, and it is for these such samples that SC performance is degraded the most.
-As the training targets never exceed $(1-\alpha)+\alpha/K$, this inaccuracy in predicting the smoothed targets could be due to epistemic uncertainty.
+However, we cannot assume that the model is equally well fit on all regions of the data distribution.
+[^3] For the sake of simplicity we avoid the term epistemic uncertainty (see Sec.
+E.2 for brief discussion).
+Through analysing the logit gradients, we provide a potential explanation: LS suppresses the max logit differently depending on how well fit the model is to a given sample – suppression is stronger(weaker) the more likely a model is correct(wrong), harming the model’s ability to separate ✓ vs ✗.
 4.1 Comparing logit gradients between CE and LS
-To better understand how LS could lead to the poor SC performance exhibited in the RC curves of Figs.
-1 and 4, we consider how LS affects logit-level training gradients.
+To better understand how LS could lead to degraded SC, we consider how LS affects logit-level training gradients.
 These are the first term in the chain rule for backpropagation and so directly contribute to all parameter gradients during training.
 We take the gradient of $\mathcal{L}^{\text{true}}$ (Eqs.
-4 and 9), which the empirical loss approximates, for a single sample,
+4 and 8), [^4] Here we assume that the empirical loss $\mathcal{L}$ (Eqs.
+2 and 6) approximates $\mathcal{L}^{\text{true}}$, in order to relate our discussion to $P_{\text{error}}$.
+However, the same analysis on the empirical loss leads to similar conclusions (see Appendix C).
 $\frac{\partial\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=-\left[\bar{\pi}_{k}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\bar{\pi}_{k}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (13)
-where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience. We can then define the regularisation gradient on the logits,
+for a single sample, where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience.
+We can then define the regularisation gradient on the logits,
 $\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (14)
 which is the difference between the LS and CE gradients.
-This intuitively represents how LS regularises training at the logit level in comparison to CE.
-Crucially, it only depends on the true distribution $\bar{\boldsymbol{\pi}}$.
+This represents how LS influences training at the logit level in comparison to CE.
+Notably, it only depends on the target $\bar{\boldsymbol{\pi}}$.
 Gradient descent involves updating weights in the opposite direction to the gradient.
-Hence $\alpha\bar{\pi}_{k}$ penalises $v_{k}$ when the true probability $\bar{\pi}_{k}$ is higher.
-The second term $-\alpha/K$ uniformly increases the logits.
-This does not affect the softmax output as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$ [44].
-3.3 Label Smoothing Exacerbates Overconfidence and Underconfidence
-Figure 7: Illustration of how the regularisation gradient (Eq.
-15), which is the difference between LS and CE gradients, affects the logits differently.
-When $U$ is lower (sharper softmax), the regularisation on the max logit is lower when epistemic uncertainty has caused overconfidence.
-When $U$ is higher (flatter softmax), the regularisation is higher when epistemic uncertainty has caused underconfidence.
-Thus, LS exacerbates both of these issues, degrading the rank ordering of $U$ and diminishing the ability of the softmax to differentiate correct  from incorrect  .
-We now consider how the regularisation gradient affects the maximum logit,
+Hence $\alpha\bar{\pi}_{k}$ suppresses $v_{k}$ when the true probability $\bar{\pi}_{k}$ is higher.
+The second term $-\alpha/K$ uniformly increases the logits for all samples, which does not affect the softmax as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$.
+4.2 Imbalanced regularisation degrades uncertainty ranking of  vs
+Figure 4: How the regularisation gradient (Eq.
+14), which is the difference between LS and CE gradients, affects the logits differently.
+LS affects the max logit differently depending on how well fit the model is for a given sample $\boldsymbol{x}$.
+On the left when $U$ is lower (sharper softmax), the regularisation on the max logit is lower when the model is poorly fit and likely to be wrong.
+On the right when $U$ is higher (flatter softmax), the regularisation is higher when the model is poorly fit and more likely to be correct.
+Thus, LS degrades the softmax’s ability to separate  vs  , hurting SC.
+We now consider how the regularisation gradient affects the maximum logit $v_{\text{max}}$,
 $\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (15)
-[…]
-To further build an intuition, we consider overconfidence and underconfidence as a result of epistemic uncertainty.
+which shows that the regularisation on $v_{\text{max}}$ decreases as the probability of error increases.
+This directly impacts softmax-based $U$ such as MSP (see Eq.
+1), especially for lower uncertainties where $v_{\text{max}}$ dominates the softmax.
+Eq.
+15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ regularised less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ regularised more.
+Thus, softmax-based $U$ will have the relative ranking of correct ✓ and incorrect ✗ predictions degraded – less uncertain on ✗ , more uncertain on ✓.
+To further build an intuition, we consider how a model may not be uniformly well-fit over the data distribution during training.
 As illustrated in Fig.
-7, when $U$ is low, and a model is overconfident due to epistemic uncertainty (i.e.
-$P_{\text{error}}$ is high), the regularisation gradient is much weaker compared to when $P_{\text{error}}$ is low for the same $U$.
-On the other hand, when $U$ is high and a model is underconfident due to epistemic uncertainty (i.e.
-$P_{\text{error}}$ is low), the regularisation gradient is much stronger compared to when $P_{\text{error}}$ is high for the same $U$.
-Thus, both overconfidence and underconfidence are exacerbated by label smoothing, degrading the model’s ability to perform selective classification.
-To test our analysis, we artificially inject epistemic uncertainty into our ImageNet evaluation using ImageNet-Sketch [76], a dataset containing sketches of each ImageNet class (see Fig.
-3 for an example).
-We combine the 50,889 ImageNet-Sketch images with the 50,000 ImageNet evaluation images and plot the new risk vs coverage for all 100,889 samples in Fig.
-8.
-We find that even though the regularisation of LS still improves the error rate, the degradation in SC from LS is exacerbated at lower coverages, where the CE-trained model maintains much lower risk.
+4, varying levels of fit with regards to the true distribution $\bar{\pi}$ leads to differing $P_{\text{error}}$ for predictions with similar $U$, resulting in the degradation described above.
+Finally, as the regularisation gradient in Eq.
+15 is proportional to $\alpha$, this explains why stronger LS leads to greater degradation.
+In summary: for a given predicted $U$ the $P_{\text{error}}$ will vary depending on how well fit the model is.
+LS suppresses the max logit more(less) when $P_{\text{error}}$ is lower(higher), degrading the ranking of  vs  for softmax-based $U$, hurting SC.
+We note that although our analysis is performed on training gradients, all our experiments are performed on evaluation data, suggesting that effects indeed generalise to unseen data.
+We also note that our explanation is purely based on the loss, and thus generalises across network architectures and tasks, matching the empirical results in Sec.
+3.
+LS empirically leads to higher $v_{\text{max}}$ on misclassifications .
+To further validate the effects of Eq.
+15, we plot the mean±$\pm$±std.
+of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ResNet-50 on evaluation data in Fig.
+5.
+We see that for LS the distribution of correct ✓  is below incorrect ✗, whilst for CE they are roughly similar.
+This provides further empirical support for Eq.
+15 which states that when using LS, $v_{\text{max}}$ is more strongly suppressed for lower $P_{\text{error}}$.
+For results on other models see Sec.
+A.2.
+LS empirically leads to increasing overconfidence on less-well-fit data.
+To further investigate how LS affects uncertainty differently depending on how well-fit/knowledgeable the model is about data, we artificially introduce distribution-shifted data that we expect our models to be worse fit on.
+We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 50,000 ImageNet evaluation images in Fig.
+6.
+Even though the regularisation of LS improves the error rate, the degradation of SC at lower coverages from LS is exacerbated by the distribution-shifted data.
 At 10% coverage of the combined evaluation set, LS leads to accepting many more errors compared to CE, with an increasing proportion originating from ImageNet-Sketch.
-This demonstrates that LS exacerbates overconfidence arising from epistemic uncertainty, reducing a model’s ability to reject errors.
+That is to say, LS leads to increasing overconfidence on less-well-fit data (ImageNet-Sketch).
+This empirically highlights the situation illustrated on the left of Fig.
+4 where confident and well-fit samples have their max logit suppressed, whilst confident but poorly fit samples do not, leading to samples with higher $P_{\text{error}}$ being relatively less uncertain (overconfident).
+For results on other models see Sec.
+A.2.
+Figure 5: Distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions on evaluation data.
+$v_{\text{max}}$ is lower for  for the LS model, whilst the distributions are roughly similar for CE.
+This empirically matches the imbalanced max logit regularisation described in Eq.
+15.
+We calculate the mean±$\pm$±std.
+in a 0.05-wide sliding window.
 @10% coverage of ImageNet + Sketch
-ImageNet
-Sketch
-CE
-#samples
-8994
-1094
-#errors
-93
-80
-error rate
-1.0
-7.3
-LS $\alpha=0.1$
-#samples
-8561
-1527
-#errors
-187
-211
-error rate
-2.2
-13.8
-LS $\alpha=0.2$
-#samples
-8171
-1917
-#errors
-244
-365
-error rate
-3.0
-19.0
-LS $\alpha=0.3$
-#samples
-7915
-2173
-#errors
-297
-541
-error rate
-3.8
-24.9
-Figure 8: Left: RC curve for ResNet-50 when evaluation is on the combination of ImageNet and ImageNet-sketch (ViT-S-16 can be found in Appendix 0.A).
-We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when epistemic uncertainty is artificially introduced.
+ImageNet Sketch
+CE #samples 8994 1094
+#errors 93 80
+error rate 1.0 7.3
+LS $\alpha=0.1$samples 8561 1527
+#errors 187 211
+error rate 2.2 13.8
+LS $\alpha=0.2$samples 8171 1917
+#errors 244 365
+error rate 3.0 19.0
+LS $\alpha=0.3$samples 7915 2173
+#errors 297 541
+error rate 3.8 24.9
+Figure 6: Left: Evaluation on the combination of ImageNet + ImageNet-sketch.
+We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced (vs Fig.
+3).
 Right: statistics @10% coverage of the combined evaluation set.
-As the level of LS $\alpha$ increases, the number of errors increases, especially the number of errors from ImageNet-Sketch.
-This shows that LS exacerbates model overconfidence arising from epistemic uncertainty.
+As the level of LS $\alpha$ increases, the number of accepted errors increases, especially from ImageNet-Sketch.
+This shows that LS leads to increasing overconfidence on less-well-fit data.
 5 Logit normalisation improves the SC of LS-trained models
 Ideally, we would like to find a way to recover from the degradation caused by LS.
-A recent study by Cattelan and Silva [4] has shown that logit normalisation can improve the SC performance of many pre-trained models.
-The logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit
-$U=-v^{\prime}_{\text{max}},\quad\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}}{\left\lVert\boldsymbol{v}\right\rVert_{p}}=\boldsymbol{v}/\bigl{(}\sum_{k}v_{k}^{p}\bigr{)}^{\frac{1}{p}}~{},$ (17)
-where $p$ is found via grid search on a validation set.
-We investigate the efficacy of this approach, applying it to our LS-trained models.
-Fig.
-9 (left) shows the SC performance with and without logit normalisation, and we indeed find that applying logit normalisation greatly improves SC performance for models trained using LS, allowing for improved error rate (@100% coverage) without compromising on SC.
-This is further visualised for semantic segmentation in Fig.
-10, where logit normalisation successfully mitigates the previously observed overconfidence caused by LS.
-However, we also find that logit normalisation does not notably improve the ability of CE-trained models.
-This aligns with the results reported [4] where certain models do not benefit from logit normalisation so the authors suggest “falling back” to the MSP.
-Figure 9: Left: RC curves showing the effect of logit normalisation.
-Logit normalisation improves SC performance on LS-trained models, but has little effect on CE-trained models.
-DeepLabV3+ can be found in Appendix 0.A.
-Right: distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately.
-We see that $v_{\text{max}}$ is higher for  for the LS-trained model, whilst the distributions are roughly the same for CE.
-This matches the max logit regularisation described in Eq.
-16.
-We calculate the mean±$\pm$±std.
-in a 0.05-wide sliding window.
-Figure 10: Visualisation of the effect of logit normalisation for semantic segmentation (same scene as Fig.
-6).
+A recent empirical study by Cattelan & Silva (2024) has shown that logit normalisation can improve the SC performance of many (but not all) pre-trained models.
+During deployment the logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit,
+$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=\boldsymbol{v}/\left\lVert\boldsymbol{v}\right\rVert_{p}=\boldsymbol{v}\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
+[…]
+Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. A.2.
+Figure 8: Visualisation of the effect of logit normalisation for segmentation (same scene as Fig.
+1).
 Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model.
+We provide more figures in Appendix F.
 5.1 Explaining the effectiveness of logit normalisation.
-Although Cattelan and Silva [4] validate this approach on a large number of pre-trained models, they provide little explanation as to why it is so effective (or indeed why it isn’t effective sometimes).
-Let us highlight the interaction between logit normalisation and label smoothing.
+Although Cattelan & Silva (2024) empirically validate this approach on a large number of pre-trained models, it remains unclear as to why it is so effective, or why it isn’t effective sometimes (further discussion in Sec.
+E.5).
+Let us shed light on the interaction between logit normalisation and LS.
 If we examine Eq.
-17, we see that it resembles the softmax.
-However, in this case, the logits are raised to power $p$ rather than being exponentiated.
+16, we see that it resembles the softmax, however, the logits are raised to power $p$ rather than exponentiated.
 Crucially, whilst the softmax is invariant to uniform shifts in the logits,
@@
 this is not the case for $\boldsymbol{v}^{\prime}$.
-In fact, we have the following inequality for positive [^5] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits.
-We include additional discussion in Sec.
-0.D.1.
+In fact, we have the following inequality for positive [^6] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits, meaning the behaviour discussed still occurs empirically (Secs.
+E.3 and 20).
 logits $\boldsymbol{v}$:
 Result 1.
-For all strictly positive vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$, containing at least two different values, and $p\in[1,+\infty[$, the ratio of the $\infty$-norm and the $p$-norm strictly decreases when summing $v$ with any uniform vector $\eta\boldsymbol{1}$, $\eta$ strictly positive:
-$\frac{||\boldsymbol{v}||_{\infty}}{||\boldsymbol{v}||_{p}}>\frac{||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}}{||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}}~{}.$ (19)
-The proof can be found in Appendix 0.C.
-Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of $v^{\prime}_{\text{max}}$.
-That is to say logit normalisation increases the estimated uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$.
-Recall that in Sec.
-3.3 we found that the maximum logit $v_{\text{max}}$ is less regularised the more likely a prediction is wrong ✗, i.e.
-when $P_{\text{error}}$ is high (Eq.
-16).
-We plot $v_{\text{max}}$ given the MSP, separately for correct ✓ and incorrect ✗ predictions in Fig.
-9, and see an interesting behaviour emerge.
-For the model trained with LS, the imbalanced logit regularisation described in Eq.
-16 has caused $v_{\text{max}}$ to become higher for misclassifications ✗ than correct predictions ✓, i.e.
-information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$.
-As Result 1 says that for the same softmax output $\boldsymbol{\pi}$, $v^{\prime}_{\text{max}}$ will decrease as $v_{\text{max}}$ increases, logit normalisation will increase uncertainty when $v_{\text{max}}$ is higher, effectively reversing the effect of the imbalanced logit regularisation.
-This will then improve the rank ordering of the uncertainty score $U$ and thus improve the SC performance of the LS-trained model.
-Given our analysis and empirical results, we strongly recommend the use of logit normalisation for LS-trained models for performing selective classification.
+For all vectors $\boldsymbol{v}\in\left(\mathbb{R}_{>0}\right)^{K}$ with at least two different values and $p\in[1,+\infty[$, the ratio of the $\infty$-norm and $p$-norm strictly decreases when summing $\boldsymbol{v}$ with any uniform vector $\eta\boldsymbol{1}$, $\eta>0$:
+$v^{\prime}_{\text{max}}(\boldsymbol{v})=||\boldsymbol{v}||_{\infty}\big{/}||\boldsymbol{v}||_{p}\quad>\quad||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}\big{/}||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}=v^{\prime}_{\text{max}}(\boldsymbol{v}+\eta\boldsymbol{1})~{}.$ (18)
+The proof can be found in Appendix D.
+Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of the normalised max logit $v^{\prime}_{\text{max}}$.
+That is to say logit normalisation increases uncertainty when the max logit $v_{\text{max}}$ is higher for the same softmax probabilities $\boldsymbol{\pi}$.
+Recall how in Sec.
+4.2 we show how imbalanced logit regularisation leads to degradation in SC.
+In particular Fig.
+5 shows how LS leads to higher $v_{\text{max}}$ given $\pi_{\text{max}}$ for errors ✗ – this implies that, independent of the value of $\pi_{\text{max}}$, information about $P_{\text{error}}$ has been encoded in $v_{\text{max}}$.
+We see now that logit normalisation will increase the uncertainty (ranking) of errors ✗  relative to correct predictions ✓  using the information in $v_{\text{max}}$, leading to improved SC for an LS model.
+That is to say logit normalisation effectively reverses the effect of the imbalanced logit regularisation from LS, improving SC.
+We note that Cattelan & Silva (2024) find that models that are less confident (on both ✓ and ✗) tend to benefit more from logit normalisation, again pointing towards LS.
+Given our analysis and empirical results, we strongly recommend logit normalisation for LS-trained models when performing SC.
 On the other hand, Fig.
-9 (right) also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE-trained model.
-This explains why, for the models trained using CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.
+5 also shows that the distributions of $v_{\text{max}}$ given $\pi_{\text{max}}$ for correct ✓ and incorrect ✗ predictions are very similar for the CE model.
+This explains why for CE, logit-normalisation does not seem to help, as $v_{\text{max}}$ does not provide useful information about $P_{\text{error}}$ given $\pi_{\text{max}}$.
 6 Related work
-5.0.1 Prediction with Rejection.
+Prediction with rejection.
 Selective classification falls into the broader problem setting of prediction with rejection.
-In the case of SC, misclassifications are to be rejected [14].
-The baseline approach is to use the MSP [30, 22] and there have been a number of proposed training [57, 33, 89] and architectural [23, 11] enhancements, however, recently the effectiveness of some of these enhancements has been called into question [15].
-In this work, we focus on only the MSP.
-Another scenario involving prediction with rejection is out-of-distribution (OOD) detection [84, 30], where data from outside of the training distribution are to be rejected.
-There is a plethora of varied research in this field [82, 68, 49, 88, 50, 34, 77, 29] extending to semantic segmentation [26, 46, 2, 24, 18].
-Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data [36, 81, 38, 5].
-Notably, Deep Ensembles [43] have arisen as a reliable method to improve performance in all three scenarios [38, 83, 45].
+In the case of SC, misclassifications are to be rejected (El-Yaniv & Wiener, 2010).
+The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023).
+Another scenario is out-of-distribution (OOD) detection (Yang et al., 2021; Hendrycks & Gimpel, 2017), where data from outside of the training distribution are to be rejected.
+There is a plethora of research in this field (Xia & Bouganis, 2022a; Sun et al., 2021; Liu et al., 2020; Zhang et al., 2023; Liu et al., 2023; Wang et al., 2022; Hendrycks et al., 2022).
+Recently, a combination of SC and OOD detection has been proposed, where the aim is to reject both misclassifications and OOD data (Jaeger et al., 2023; Xia & Bouganis, 2022b; Kim et al., 2023).
+Notably, Deep Ensembles (Lakshminarayanan et al., 2020) have arisen as a reliable method to improve performance in all three scenarios (Kim et al., 2023; Xia & Bouganis, 2023; Laurent et al., 2023; 2024).
 We believe, given the results of this work, that extending the investigation of LS (and other training enhancements) to other scenarios involving prediction with rejection is an important avenue of future work.
-5.0.2 Mixup.
-Mixup [87] and its variants [86, 17, 62, 54, 3] are a set of regularisation techniques that involve interpolating between random pairs of samples at training time, modifying both inputs and targets.
-They are commonly found in many ImageNet training recipes [73, 52, 51] and often used in conjunction with label smoothing [80, 74, 53].
-We remark that research into how Mixup and its variants affect SC is a particularly salient avenue of future work as it similarly “softens” the classification training targets.
-5.0.3 Label Smoothing.
-Beyond prediction with rejection, label smoothing [70] is well investigated.
-It has been shown to improve model calibration [59, 58, 47, 10] as well as accuracy when training under label noise [56].
-Knowledge distillation [31], where soft labels are provided by a teacher network, is also commonly linked and combined with LS [59, 21, 65, 6, 85] due to their similarity.
-Interestingly, in a similar vein to our work, pre-training using LS has been shown to harm transfer learning [40].
+Mixup.
+Mixup (Zhang et al., 2018) and its variants (Yun et al., 2019; Franchi et al., 2021; Pinto et al., 2022; Liu et al., 2022d; Bouniot et al., 2023) are a set of regularisation techniques that involve interpolating between random pairs of samples at training time, modifying both inputs and targets.
+They are commonly found in many ImageNet training recipes (Tan & Le, 2021; Liu et al., 2021; 2022b) and are often used in conjunction with label smoothing (Wightman et al., 2021; Touvron et al., 2021; Liu et al., 2022c).
+Research into how Mixup affects SC is a particularly salient avenue of future work as (Zhu et al., 2024) also observe empirically that it can have a negative impact.
+Label smoothing.
+Beyond prediction with rejection, LS is well-explored across various contexts.
+It has been shown to improve model calibration (Müller et al., 2019; Chun et al., 2020; Mukhoti et al., 2020; Liu et al., 2022a) as well as accuracy when training under label noise (Lukasik et al., 2020).
+Knowledge distillation (Hinton et al., 2015), where soft labels are provided by a teacher network, is also commonly linked and combined with LS (Müller et al., 2019; Gao et al., 2020; Yuan et al., 2020; Shen et al., 2021; Chandrasegaran et al., 2022) due to their similarity.
+Interestingly, in a similar vein to our work, pre-training using LS has been shown to harm transfer learning (Kornblith et al., 2021).
 7 Concluding remarks
-In this work, we investigate the effect that label smoothing has on a model’s ability to perform selective classification.
-Our experiments across a range of architectures and tasks reveal that the use of label smoothing leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves classification performance.
-By analysing the logit-level gradients of the label smoothing loss, we are able to explain this behaviour – label smoothing increasingly regularises the max logit as the true probability of error decreases, exacerbating overconfidence and underconfidence.
-We then investigate logit normalisation as a potential method to improve the degraded selective performance of label-smoothing-trained models.
-We find it to be highly effective, as it, in essence, reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher.
+In this work, we elucidate the effect of label smoothing (LS) on selective classification (SC).
+Our experiments across various tasks and architectures show that LS leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves accuracy.
+By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS regularises/suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions.
+We then investigate post-hoc logit normalisation as a method to improve the degraded SC performance caused by LS.
+We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher.
 We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
-6.0.1 Acknowledgments.
-This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-[AD011011970R3]). GX’s PhD is funded jointly by Arm and the EPSRC.
-Supplementary Material
-Appendix 0.A Additional Results
+Reproducibility statement
+To ensure transparency, we use publicly available datasets: CIFAR-100, ImageNet, and Cityscapes.
+Our detailed experimental methods are outlined in Appendix B.
+Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix D.
+To help replicate our work, we will soon share our source code on GitHub, including configuration files and the training code. We also plan to release our most important models on Hugging Face.
+Ethics
+Our primary objective is to contribute to enhancing the reliability of machine learning methods, with a particular focus on raising awareness about the negative effects of label-smoothing on selective classification.
+While our work aims to improve the robustness of ML systems, we acknowledge the potential risk that these advancements could be misapplied in harmful ways.
+Appendix A Additional results
+A.1 Illustrative examples of ImageNet overconfidence
+Fig. 9 shows a few examples of overconfident misclassifications on the ImageNet evaluation data by our LS-trained ResNet-50.
+Figure 9: Illustrative examples of overconfident errors performed by our LS-trained ResNet-50 on evaluation data.
+Even though $P_{\text{error}}$ is high in all cases (e.g.
+due to multiple possible labels), the model predicts low (ranking) uncertainty.
+Note that even though the CE model is wrong as well, it has assigned higher ranking uncertainties, reflecting its superior SC ability shown in Fig.
+3.
+A.2 Complete results for ViT and DeepLabV3+
 We include additional experimental results that mirror those found in the main paper on ResNet-50, for ViT-S-16 and DeepLabV3+:
-  • Fig. 11 shows ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch. The behaviour mirrors that of ResNet-50 in the main paper.
-  • Fig. 12 shows the effectiveness of logit normalisation on DeepLabV3+ (ResNet-101) on Cityscapes.
+  • Fig.
+10 and Tab.
+1 show full ResNet-50 and ViT-S-16 results on ImageNet + ImageNet-Sketch.
+We see that the behaviour of ViT-S-16 is similar to ResNet-50 (increasing numbers of confident errors from sketch as $\alpha$ increases), but less pronounced.
   • Figs.
-13 and 14 shows the distribution of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ViT-S-16 and DeepLabV3+.
+11 and 12 shows the distribution of $v_{\text{max}}$ given $\pi_{\text{max}}$ for ViT-S-16 and DeepLabV3+.
 We see that similarly to ResNet-50, for LS, the distribution of $v_{\text{max}}$ is higher for errors ✗.
 Although it is less obvious, it is clear for both ViT-S-16 and DeepLabV3+ that for higher $\pi_{\text{max}}$ the standard deviations overlap much less than for CE.
-@10% coverage of ImageNet + Sketch
-ImageNet
-Sketch
-CE
-#samples
-9528
-560
-#errors
-88
-43
-error rate
-0.9
-7.7
-LS $\alpha=0.1$
-#samples
-9332
-756
-#errors
-130
-83
-error rate
-1.4
-11.0
-LS $\alpha=0.2$
-#samples
-9117
-971
-#errors
-177
-164
-error rate
-1.9
-16.9
-LS $\alpha=0.3$
-#samples
-8966
-1122
-#errors
-208
-195
-error rate
-2.3
-17.4
-Figure 11: ViT-S-16 evaluated on the combination of ImageNet and ImageNet-Sketch.
-The results mirror ResNet-50 from the main paper.
-LS leads to increased overconfidence when epistemic uncertainty is introduced, as shown through the increase in the failure rate amongst high-confidence samples.
-Figure 12: RC curve showing the effect of logit normalisation for DeepLabV3+ (ResNet-101).
-The behaviour is similar to the results in the main paper.
-Logit normalisation is effective in improving the performance of the LS-trained model, however, makes little difference to the CE-trained model.
-Figure 13: The max logit $v_{\text{max}}$ as a function of the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16.
-Similarly to the main paper, the distribution is higher for errors  for the LS-trained model.
+  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix F.
+Figure 10: RC curves when evaluating on the combination of ImageNet and ImageNet-sketch.
+We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced.
+This shows that LS leads to increasing overconfidence on less-well-fit data.
+Table 1: Statistics @10% coverage of the combined evaluation set. As the level of LS $\alpha$ increases, the number of errors increases, especially the number of errors from ImageNet-Sketch.
+ResNet-50 @10% coverage of ImageNet + Sketch
+ImageNet Sketch
+CE #samples 8994 1094
+#errors 93 80
+error rate 1.0 7.3
+LS $\alpha=0.1$samples 8561 1527
+#errors 187 211
+error rate 2.2 13.8
+LS $\alpha=0.2$samples 8171 1917
+#errors 244 365
+error rate 3.0 19.0
+LS $\alpha=0.3$samples 7915 2173
+#errors 297 541
+error rate 3.8 24.9
+ViT-S-16 @10% coverage of ImageNet + Sketch
+ImageNet Sketch
+CE #samples 9528 560
+#errors 88 43
+error rate 0.9 7.7
+LS $\alpha=0.1$samples 9332 756
+#errors 130 83
+error rate 1.4 11.0
+LS $\alpha=0.2$samples 9117 971
+#errors 177 164
+error rate 1.9 16.9
+LS $\alpha=0.3$samples 8966 1122
+#errors 208 195
+error rate 2.3 17.4
+Figure 11: Distribution of the max logit $v_{\text{max}}$ given MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16.
+Similarly to the main paper, the distribution of errors  is higher than correct predictions  for the LS-trained model.
 We calculate the mean±$\pm$±std.
@@
 Figure 12: The same as Fig. 11 but for DeepLabV3+ (ResNet-101)
-0.A.1 Small-Scale CIFAR Experiments
-We also include experimental results on small-scale $32\times 32$ CIFAR-100 [41].
-We train a DenseNet-BC [32] ($k=12,L=100$) to show further generality over model architecture families.
+Figure 13: RC curve showing the effect of logit normalisation for ViT-S-16 and DeepLabV3+ (ResNet-101).
+The behaviour is similar to the results in the main paper.
+Logit normalisation is effective in improving the performance of the LS-trained models, bringing them close to the CE models.
+However, logit normalisation makes little difference to the CE-trained model.
+A.3 Small-scale CIFAR experiments
+We also include experimental results on small-scale $32\times 32$ CIFAR-100 (Krizhevsky, 2009).
+We train a DenseNet-BC (Huang et al., 2017) ($k=12,L=100$) to show further generality over model architecture families.
 Figs.
-15, 16 and 17 show results that mirror those found in the main paper for ResNet-50 on ImageNet, although we note that LS does not improve top-1 error rate in this case.
+14, 15 and 16 show results that mirror those found in the main paper for ResNet-50 on ImageNet, although we note that LS does not improve top-1 error rate in this case.
 Figure 14: RC curves for DenseNet on CIFAR-100 – LS degrades SC
@@
 Figure 16: DenseNet on CIFAR-100 – LS leads to higher $v_{\text{max}}$ on errors
-Appendix 0.B Reproducibility
-Alongside this document, we provide a code demo to train two ResNet-20 [27] on CIFAR-10 [41] with cross-entropy and label smoothing and compare the corresponding Risk-Coverage curves.
-We recall that all the code used in this project will be made available shortly to ensure exact reproducibility.
-We also plan to release our most important models on Hugging Face.
-Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix 0.A.
+Appendix B Reproducibility
+We recall that all the code used in this project will be made available soon to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
+Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix A.
 All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
-0.B.1 Image Classification
+B.1 Image classification
 DenseNet – CIFAR100.
 For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set.
-We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov [60, 69], a momentum of 0.9, and $1\times 10^{-4}$ weight decay.
+We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov (Nesterov, 1983; Sutskever et al., 2013), a momentum of 0.9, and $1\times 10^{-4}$ weight decay.
 We divide the learning rate by ten after 150 and 225 epochs.
@@
 We do not perform model selection and keep the last checkpoint.
-ViT-16-S – ImageNet.
-For our ViT-16-S trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW [55] with the $\beta$s equal to 0.9 and 0.999.
+ViT-S-16 – ImageNet.
+For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999.
 We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate.
-The models are trained with non-adaptive sharpness-aware minimisation (SAM) [16, 8] with $\rho=0.2$.
-We use a dropout [66] rate of 0.1 but no attention dropout.
+The models are trained with non-adaptive sharpness-aware minimisation (SAM) (Foret et al., 2021; Chen et al., 2022) with $\rho=0.2$.
+We use a dropout (Srivastava et al., 2014) rate of 0.1 but no attention dropout.
 We transform the training images with a standard random resized crop to 224$\times$224 pixels using bicubic interpolation and a random horizontal flip.
@@
 We use the same transformation of the images as for the ViT and select the last checkpoint for inference.
-We use the same evaluation set as for the ViT for logit normalisation.
-0.B.2 Semantic Segmentation
+We use the same validation set as for the ViT for logit normalisation.
+B.2 Semantic segmentation
 Deeplabv3+.
-We train a Deeplabv3+ on CityScapes [12] with a ResNet-101 [27] backbone pre-trained on ImageNet [63].
-We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy [48] with a power of 0.9.
+We train a Deeplabv3+ on CityScapes (Cordts et al., 2016) with a ResNet-101 (He et al., 2016) backbone pre-trained on ImageNet (Russakovsky et al., 2015).
+We use stochastic gradient descent with a base learning rate of 0.01, divided by 10 for the backbone weights, and reduced following the "poly" policy (Liu et al., 2015) with a power of 0.9.
 The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$.
@@
 We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
-Appendix 0.C Result and Proof
+B.3 Logit Normalisation
+[…]
+Appendix C Analysis of empirical loss (one-hot labels)
+We can perform a similar analysis as in Sec.
+3 using the empirical loss (Eqs.
+2 and 6) rather than $\mathcal{L}^{\text{true}}$ as in the main paper.
+We leave this to the Appendix as the conclusions are similar to analysing $\mathcal{L}^{\text{true}}$, however, the presence of one-hot targets makes it less convenient to reason about the probability of error $P_{\text{error}}$ and how well fit the model is to the true conditional distribution $P_{\text{data}}(y|\boldsymbol{x})$ for a given data sample.
+Taking the gradients of Eqs.
+2 and 6 as in Eq.
+13,
+$\frac{\partial\mathcal{L}_{\text{CE}}}{\partial v_{k}}=-\left[\delta_{y\omega_{k}}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\delta_{y\omega_{k}}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (19)
+which in turn gives the regularisation gradient,
+$\frac{\partial\mathcal{L}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
+which is once again independent of the model output $\boldsymbol{\pi}$.
+In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different regularisation gradients for the max logit $v_{\text{max}}$,
+[…]
+We can see that the max logit is more strongly regularised/suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper.
+When the model is correct during training, the max logit is suppressed, but when it is incorrect the max logit is not suppressed.
+Thus the uncertainty ranking of correct ✓ vs incorrect ✗ data samples is degraded, harming SC.
+This leads us to the same conclusions as in Sec.
+4 and also aligns with the behaviour in Fig.
+5 (right) where $v_{\text{max}}$ is higher for errors given the value of MSP.
+Appendix D Result and proof
 Result.
@@
 ∎
-Appendix 0.D Discussions
-0.D.1 On the Assumption that Logits are Positive
+Appendix E Discussions
+E.1 $U$ other than MSP
+In the main body of the paper we focus solely on MSP as our uncertainty score $U$.
+Fig.
+17 shows how LS affects the SC behaviour of MSP (top) compared two other softmax scores: DOCTOR (Granese et al., 2021) ($U=-||\boldsymbol{\pi}||_{2}$) (middle) and entropy ($U=H(\boldsymbol{\pi})=-\sum_{k}\pi_{k}\log\pi_{k}$) (bottom).
+We see that the behaviour is very similar across all three softmax scores, with increasing the level of LS $\alpha$ degrading all the scores.
+As MSP is the (marginally) best performing and the most commonly used, we thus choose to focus on it for the main paper.
+Fig.
+18 shows the SC performance of Energy ($U=-\log\sum_{k}\exp v_{k}$) (Liu et al., 2020), a popular OOD detection score, compared to MSP with and without LS.
+We see that Energy performs much worse than MSP at SC.
+This aligns with a large body of existing work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Kim et al., 2023; Yang et al., 2024; Zhu et al., 2024) that empirically finds that uncertainty scores designed for OOD detection perform poorly at detecting misclassifications and are thus not suitable for SC.
+Thus we choose not to investigate any OOD detection scores in the main paper.
+We remark that LS also has a strong negative effect on the SC performance of Energy.
+This is unsurprising as Energy is dominated by the max logit.
+An important future direction would be to investigate the effect of LS on various OOD detection scores on the task of OOD detection.
+Finally, we also choose to omit various training/architecture-based approaches such as (Ziyin et al., 2019; Geifman & El-Yaniv, 2019; Moon et al., 2020; Zhu et al., 2024), as we aim to focus our investigation solely on understanding the training effects of LS.
+We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use-cases and so may be more likely to be chosen by a practitioner.
+Besides, recent work (Feng et al., 2023) has suggested that MSP applied to the methods in (Ziyin et al., 2019; Geifman & El-Yaniv, 2019) actually performs better than the $U$s proposed in them (reject logit/selection head).
+Figure 17: The effect of LS on different softmax scores: MSP (top), DOCTOR (middle), Entropy (bottom). We see that the behaviour is very similar, with MSP being the best performer.
+Figure 18: SC performance of MSP and Energy (OOD detection score).
+Energy significantly underperforms MSP.
+This behaviour is in line with existing work that shows that uncertainty scores designed for OOD detection are not suitable for SC.
+E.2 Aleatoric and epistemic uncertainty
+We recognise that the conceptual decomposition of predictive uncertainty into aleatoric and epistemic uncertainty (Gal, 2016; Hüllermeier & Waegeman, 2021; Kirsch, 2024) can be applied to the discussion throughout this paper, and that some readers may be confused as to why we do not use these terms.
+We choose to use simpler, more direct language as we believe it more efficiently conveys our discussion, reduces the number of concepts to introduce and also reduces any potential confusion.
+E.3 On the assumption that logits are positive
 Figure 19: Mean±$\pm$±std after logits have been sorted highest to lowest for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set.
@@
 Figure 20: $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ within bins of similar MSP for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set.
-Generally, $v^{\prime}_{\text{max}}$ increases as $v_{\text{max}}$ decreases, showing empirically how logit normalisation increases uncertainty for higher $v_{\text{max}}$.
+Generally, $v^{\prime}_{\text{max}}$ decreases as $v_{\text{max}}$ increases, showing empirically how logit normalisation increases uncertainty for higher $v_{\text{max}}$.
 We calculate the mean±$\pm$±std.
@@
 However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits.
-This is intuitive as exponentiating or raising to power $p$ ($p\geq 4$ in our experiments) will amplify the larger logits.
+This is intuitive as exponentiating or raising to power $p>1$ will amplify the larger logits.
 This is shown in Fig.
-18, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set.
+19, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case).
 Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive.
-As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to penalise higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
-We can consider the scenario where we add $\eta$ to only the top-$k$ logits.
+As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
+We can also consider the scenario where we add $\eta$ to only the top-$k$ logits.
 This more aptly describes the empirical logit behaviour compared to adding $\eta$ to all logits, as the values of the lower ranking logits vary much less than the higher ranking ones (Fig.
-18).
+19).
 Here we would expect $\pi_{\text{max}}(\boldsymbol{v})$ to increase very slightly, but would expect $v^{\prime}_{\text{max}}(\boldsymbol{v})$ to decrease as the numerator of Eq.
@@
 We plot the mean±$\pm$±std of $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ for samples in different MSP bins.
-We see clearly that in almost all cases, for samples with similar $\pi_{\text{max}}$, the normalised max logit $v^{\prime}_{\text{max}}$ increases as the original max logit $v_{\text{max}}$ decreases.
-As such, logit normalisation is able to improve the SC performance of LS-trained models.
-0.D.2 Existing Benchmarks and Training Recipes
-Although we do not exhaustively search all training recipes for all models benchmarked in [20, 4], we do provide a number of examples of evaluated models trained with label smoothing.
+We see clearly that in almost all cases, for samples with similar $\pi_{\text{max}}$, the normalised max logit $v^{\prime}_{\text{max}}$ decreases as the original max logit $v_{\text{max}}$ increases.
+As shown in Figs.
+5, 16, 11 and 12, LS leads to misclassifications ✗ having higher $v_{\text{max}}$ than correct predictions ✓.
+Thus, logit normalisation is able to improve the SC performance of LS-trained models by penalising the confidence of higher max logit values.
+E.4 Existing benchmarks and training recipes with LS
+Although we do not exhaustively search all training recipes for all models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024), we do provide a number of examples of evaluated models trained with label smoothing.
 We also provide links to publicly available training repositories, as not all papers mention label smoothing even when it is used in training.
-Upon inspection of [20, 4], these models do in fact seem to underperform at selective classification (and Cattelan and Silva [4] report that their AURCs benefit from logit normalisation).
+Upon inspection of (Galil et al., 2023; Cattelan & Silva, 2024), these models do in fact seem to underperform at selective classification (and Cattelan & Silva (2024) report that their AURCs benefit from logit normalisation).
 Galil et al.
-[20] state that some of their best performing ViT models [13, 67, 8] are trained with label smoothing (their Tab.1).
+(2023) state that some of their best performing (at SC) ViT models (Dosovitskiy et al., 2021; Steiner et al., 2022; Chen et al., 2022) are trained with label smoothing (their Tab.1).
 However, after inspecting both the original papers and open-source repositories [^7] https://github.com/google-research/vision_transformer of the aforementioned work we were unable to find any confirmation of the use of label smoothing.
+E.5 Logit normalisation in (Cattelan & Silva, 2024)
+Explanations of effectiveness.
+The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways.
+However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B.
+They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix D).
+We believe their explanation is ultimately incomplete as:
+  1. 1.
+    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification.
+This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties.
+It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples Zhu et al.
+(2024).
+  2. 2.
+    Although they elucidate some of the mechanics of logit normalisation, showing that logit normalisation reduces the impact of smaller logits on uncertainty, they do not link it to any observed model behaviour that differs between correct ✓ vs incorrect ✗ predictions, instead just speculating that models may be underconfident on certain sets of samples.
+To explain why an approach improves SC, we need to know how it treats ✓ and ✗ differently to improve the rank ordering of uncertainties between ✓ and ✗.
+On the other hand, our explanation focuses solely on the relative ranking of uncertainties between ✓ vs ✗ samples and is able to link the mechanics of logit normalisation to how the behaviour of correct ✓ vs incorrect ✗ predictions differ under label smoothing (Figs.
+5 and 1).
+Logit centralisation.
+Cattelan & Silva (2024) suggest an additional step of logit centralisation,
+$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k},$ (28)
+which in practice brings most of the lower valued logits close to zero if they are not already.
+They note that for the majority of models this step is unnecessary, and we also find it unnecessary in our experiments so we omit it for the sake of simplicity.
+We note that according to the analysis in Sec.
+5.1 and Sec.
+E.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits, which centralisation seems to empirically ensure by making most of the smallest logits close to zero (Cattelan & Silva, 2024).
+Appendix F Additional segmentation results
+In this section, we complete the picture by providing additional semantic segmentation results to Figures 1 and 8.
+As in the main paper, the following predictions are performed by a ResNet-101-based DeepLab-v3+ trained on Cityscapes.
+In these figures, we provide the rank of all the predictions – not only errors – to show the difference between the model’s behavior on its errors but also correct predictions.
+The rank of the correct predictions ✓ is slightly greyed out compared to the rank of the errors ✗.
+Figure 21 presents the full ranks corresponding to the scene used in the main paper.
+We will provide other examples to the readers upon request.
+Later, we will make our models available to ease the reproducibility of our work.
+Figure 21: We provide all the predicted ranks on our segmentation plots to understand where the model predicted higher relative confidence.
+The pixel uncertainty rankings provided by the model trained with CE (left) are smoother than those of the model trained with LS $\alpha=0.2$ (right).
+Figure 22 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are smoother than those of the label-smoothing-based one.
+The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly decreases nearby the boundary of objects.
+However, the boundaries are sometimes wrongly predicted, leading to (rank-wise) highly confident errors ✗.
+The pixels corresponding to the boundaries of objects, where naturally the probability of error is higher and which are most sensitive to label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model.
+This explains the high-confidence errors ✗ in the misclassified boundaries of several objects, either in Figs.
+1 and 22, and aligns with the imbalanced regularisation of Eq.
+15.
+We also provide Figure 23 to show that the logit normalisation method is not always successful in solving the ranking of the predictions.
+In this figure, we see that a large part of the misclassified pixels (lower left) keeps low uncertainty ranks either using the CE-based model or the LS-based model while using logit normalisation.
+However, there are some improvements on the label-smoothing side, where the original low-ranked boundary errors ✗ are less (rankwise) confident, such as the roof of the car and the right border of the median strip.
+Interestingly, LS-based high confidence errors in Figure 23 mainly correspond to a change of texture of the median strip – likely associated with an uncertainty on the true label – and correlate with those of the CE-based model.
+For brevity, we provide an image for a logit normalisation with norm dimension $p=5$, but the results do not differ much from the other values.
+As stated above, we will provide additional figures to readers upon request.
+Figure 22: Logit normalisation improves on boundary-related very confident errors ✗, but not on the bottom right errors ✗.
+Figure 23: Logit normalisation does not always fix selective classification when used on a model trained with a label-smoothing (right).
+However, boundary errors ✗ have lower (ranked) confidence, such as for the car’s roof.

### v3

@@
 In this work, we first demonstrate empirically across an extended range of large-scale tasks and architectures that LS consistently degrades SC.
-We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by regularising the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong.
+We then address a gap in existing knowledge, providing an explanation for this behaviour by analysing logit-level gradients: LS degrades the uncertainty rank ordering of correct vs incorrect predictions by suppressing the max logit more when a prediction is likely to be correct, and less when it is likely to be wrong.
 This elucidates previously reported experimental results where strong classifiers underperform in SC.
@@
 Furthermore, linking back to our gradient analysis, we again provide an explanation for why such normalisation is effective.
+Project page: https://ensta-u2is-ai.github.io/Understanding-Label-smoothing-Selective-classification/
 1 Introduction
@@
 ResNet-50 ImageNet MSP CE LS $\alpha=0.1$ LS $\alpha=0.2$ LS $\alpha=0.3$ LS $\alpha=0.2$ – logit norm
-%coverage@risk=1$\uparrow$ 17.69 0.2 0.53 0.49 21.9
+%coverage@risk=1$\uparrow$ 15.66 0.05 0.06 0.02 22.31
 Figure 1: Top: LS causes overconfidence for semantic segmentation.
@@
 1), even if it may improve accuracy.
-Moreover, we show that the degradation worsens with stronger LS.
+Moreover, we find that the degradation worsens with stronger LS.
 As LS can be found in the training recipes of many of the models evaluated in (Galil et al., 2023; Cattelan & Silva, 2024), this suggests LS as one potential cause for previously unexplained negative results where strong classifiers underperform at SC.
@@
     We address a gap in the understanding of LS by providing an explanation of this behaviour through analysing the logit-level gradients of the LS loss.
-We show that the amount LS regularises/suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with regularisation increasing(decreasing) the more likely a prediction is correct(wrong).
+We show that the amount LS suppresses the max logit directly corresponds to the true probability of error $P_{\text{error}}$, with suppression increasing(decreasing) the more likely a prediction is correct(wrong).
 This leads to relatively higher uncertainty on correct predictions and lower uncertainty on misclassifications, degrading the ranking of uncertainties and hurting SC compared to vanilla CE.
@@
 Moreover, we elucidate this effectiveness through the lens of our gradient-based analysis.
-Linking back to the imbalanced logit regularisation of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
+Linking back to the imbalanced logit suppression of LS, we find that logit normalisation compensates for this effect by increasing uncertainty as the max logit increases.
 2 Preliminaries
+For a glossary of notation see Appendix A.
 Consider a $K$-class neural network classifier with parameters $\boldsymbol{\theta}$ that models the conditional distribution $P(y|\boldsymbol{x};\boldsymbol{\theta})$ of labels $y\in\mathcal{Y}=\{\omega_{k}\}_{k=1}^{K}$ given inputs $\boldsymbol{x}\in\mathcal{X}=\mathbb{R}^{D}$.
@@
 $=\mathbb{E}_{p_{\text{data}}(\boldsymbol{x})}\left[\text{KL}\left[\bar{\boldsymbol{\pi}}(\boldsymbol{x})||\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})\right]\right]+\text{const.}~{}=\mathcal{L}_{\text{CE}}^{\text{true}}(\boldsymbol{\theta})~{},$ (4)
-where $\delta$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence.
+where $\delta_{ij}=1\text{ if }i=j,\text{ and }0\text{ if }i\neq j$ is the Kronecker delta and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback–Leibler divergence.
 We use $\bar{\boldsymbol{\pi}}(\boldsymbol{x})\in[0,1]^{K}$ as a shorthand for the true conditional categorical, i.e.
@@
 In line with previous work (Xia & Bouganis, 2022b; Jaeger et al., 2023; Zhu et al., 2024), we find that OOD detection (Yang et al., 2021) scores perform poorly at SC, and also omit them (see Sec.
-E.1 for discussion and additional results).
+F.1 for discussion and additional results).
 Over/Underconfidence.
@@
 Experimental details.
-We investigate large-scale [^2] We additionally provide $32\times 32$ small-scale CIFAR experiments in Sec.
-A.3.
+We investigate large-scale [^2] We additionally provide small-scale CIFAR, tabular and text experiments in Secs.
+B.3, B.4 and B.5.
 image classification on ImageNet-1k (Russakovsky et al., 2015) and semantic segmentation (pixel-level classification) on Cityscapes (Cordts et al., 2016).
-For ImageNet, we evaluate on the original validation set and randomly split 50,000 images from the training set for validation.
+For ImageNet, we randomly split the validation set into 10,000 validation and 40,000 evaluation images.
 For Cityscapes, we randomly split the original validation set into 100 validation and 400 evaluation images.
-To estimate the risk, we extract the model output prior to the final interpolation and subsample 5000 labelled pixels per image at random.
+To estimate the risk for the RC curves, we subsample 5000 labelled pixels per image at random.
 We evaluate on the same pixels between models.
@@
 For semantic segmentation on Cityscapes, we train DeepLabV3+ (Chen et al., 2018) (ResNet-101 backbone) using only random cropping, flipping and colour jitter for augmentations.
-Full training details for reproducibility can be found in the Appendix B.
-We will release our code in the future.
+Extended training details for reproducibility can be found in the Appendix C.
+We release our code on GitHub.
 3.1 Label smoothing degrades selective classification
 Figure 3: Risk-coverage plots for different levels of LS $\alpha$ for different models and tasks (ImageNet classification and Cityscapes semantic segmentation).
-Although LS may improve error rate at 100% coverage, increasing $\alpha$ increasingly and consistently degrades SC performance.
+Although it may improve error rate/accuracy at 100% coverage, label smoothing consistently degrades SC performance.
 To examine the effects of LS on SC, we plot RC curves in Fig.
@@
 We include additional discussion of existing benchmarks and model checkpoints in Sec.
-E.4.
+F.6.
 4 Towards explaining the negative effect of LS on SC
@@
 [^3] For the sake of simplicity we avoid the term epistemic uncertainty (see Sec.
-E.2 for brief discussion).
+F.2 for brief discussion).
 Through analysing the logit gradients, we provide a potential explanation: LS suppresses the max logit differently depending on how well fit the model is to a given sample – suppression is stronger(weaker) the more likely a model is correct(wrong), harming the model’s ability to separate ✓ vs ✗.
@@
 2 and 6) approximates $\mathcal{L}^{\text{true}}$, in order to relate our discussion to $P_{\text{error}}$.
-However, the same analysis on the empirical loss leads to similar conclusions (see Appendix C).
+However, the same analysis on the empirical loss leads to similar conclusions (see Appendix D).
 $\frac{\partial\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=-\left[\bar{\pi}_{k}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\bar{\pi}_{k}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (13)
 for a single sample, where in a slight abuse of notation we omit the outer expectation over $p_{\text{data}}(\boldsymbol{x})$ for convenience.
-We can then define the regularisation gradient on the logits,
-$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (14)
+We can then define the suppression gradient on the logits,
+$\frac{\partial\mathcal{L}^{\text{true}}_{\text{{\color[rgb]{0,0,0}sup}}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}^{\text{true}}-\mathcal{L}_{\text{CE}}^{\text{true}})}{\partial v_{k}}=\frac{\partial\mathcal{L}_{\text{LS}}^{\text{true}}}{\partial v_{k}}-\frac{\partial\mathcal{L}_{\text{CE}}^{\text{true}}}{\partial v_{k}}=\alpha\left[\bar{\pi}_{k}-1/K\right]=\alpha\bar{\pi}_{k}-\alpha/K~{},$ (14)
 which is the difference between the LS and CE gradients.
@@
 The second term $-\alpha/K$ uniformly increases the logits for all samples, which does not affect the softmax as it is invariant to uniform logit shifts $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$, $\eta\in\mathbb{R}$.
-4.2 Imbalanced regularisation degrades uncertainty ranking of  vs
-Figure 4: How the regularisation gradient (Eq.
-14), which is the difference between LS and CE gradients, affects the logits differently.
+4.2 Imbalanced suppression degrades uncertainty ranking of  vs
+Figure 4: How the suppression gradient (Eq.
+14), i.e.
+the difference between LS and CE gradients, affects the logits differently.
 LS affects the max logit differently depending on how well fit the model is for a given sample $\boldsymbol{x}$.
-On the left when $U$ is lower (sharper softmax), the regularisation on the max logit is lower when the model is poorly fit and likely to be wrong.
-On the right when $U$ is higher (flatter softmax), the regularisation is higher when the model is poorly fit and more likely to be correct.
+In the left two when $U$ is lower (sharper softmax), the suppression on the max logit is lower when the model is poorly fit and likely to be wrong.
+In the right two when $U$ is higher (flatter softmax), the suppression is higher when the model is poorly fit and more likely to be correct.
 Thus, LS degrades the softmax’s ability to separate  vs  , hurting SC.
-We now consider how the regularisation gradient affects the maximum logit $v_{\text{max}}$,
-$\frac{\partial\mathcal{L}^{\text{true}}_{\text{reg}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (15)
-which shows that the regularisation on $v_{\text{max}}$ decreases as the probability of error increases.
-This directly impacts softmax-based $U$ such as MSP (see Eq.
-1), especially for lower uncertainties where $v_{\text{max}}$ dominates the softmax.
+We now consider how the suppression gradient affects the maximum logit $v_{\text{max}}$,
+$\frac{\partial\mathcal{L}^{\text{true}}_{\text{{\color[rgb]{0,0,0}sup}}}}{\partial v_{\text{max}}}=\alpha\bar{\pi}_{\hat{y}}-\alpha/K=\alpha[1-P_{\text{error}}]-\alpha/K,$ (15)
+which shows that the suppression on $v_{\text{max}}$ decreases as the probability of error increases.
+This directly impacts softmax-based $U$ such as MSP (see Eqs.
+1 and 12), as the exponentiation of the softmax will be dominated by the largest logits, especially the max logit $v_{\text{max}}$.
+The max logit will be especially dominant on high-confidence samples which are the last to be rejected.
 Eq.
-15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ regularised less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ regularised more.
+15 suggests that for predictions that share the same value of $U$, those with higher $P_{\text{error}}$ (more likely ✗), will have $v_{\text{max}}$ suppressed less, whilst predictions with lower $P_{\text{error}}$ (more likely ✓) will have $v_{\text{max}}$ suppressed more.
 Thus, softmax-based $U$ will have the relative ranking of correct ✓ and incorrect ✗ predictions degraded – less uncertain on ✗ , more uncertain on ✓.
@@
 4, varying levels of fit with regards to the true distribution $\bar{\pi}$ leads to differing $P_{\text{error}}$ for predictions with similar $U$, resulting in the degradation described above.
-Finally, as the regularisation gradient in Eq.
+Finally, as the suppression gradient in Eq.
 15 is proportional to $\alpha$, this explains why stronger LS leads to greater degradation.
@@
 For results on other models see Sec.
-A.2.
+B.2.
 LS empirically leads to increasing overconfidence on less-well-fit data.
 To further investigate how LS affects uncertainty differently depending on how well-fit/knowledgeable the model is about data, we artificially introduce distribution-shifted data that we expect our models to be worse fit on.
-We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 50,000 ImageNet evaluation images in Fig.
+We use ImageNet-Sketch (Wang et al., 2019), a dataset containing sketches of each ImageNet class and evaluate on the combination of the 50,889 ImageNet-Sketch and 40,000 ImageNet evaluation images in Fig.
 6.
@@
 For results on other models see Sec.
-A.2.
+B.2.
 Figure 5: Distribution of the max logit $v_{\text{max}}$ given the MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions on evaluation data.
 $v_{\text{max}}$ is lower for  for the LS model, whilst the distributions are roughly similar for CE.
-This empirically matches the imbalanced max logit regularisation described in Eq.
+This empirically matches the imbalanced max logit suppression described in Eq.
 15.
@@
 ImageNet Sketch
-CE #samples 8994 1094
-#errors 93 80
-error rate 1.0 7.3
-LS $\alpha=0.1$samples 8561 1527
-#errors 187 211
-error rate 2.2 13.8
-LS $\alpha=0.2$samples 8171 1917
-#errors 244 365
-error rate 3.0 19.0
-LS $\alpha=0.3$samples 7915 2173
-#errors 297 541
-error rate 3.8 24.9
+CE #samples 7940 1148
+#errors 94 74
+error rate 1.2 6.4
+LS $\alpha=0.1$samples 7422 1666
+#errors 174 296
+error rate 2.3 17.8
+LS $\alpha=0.2$samples 6808 2280
+#errors 225 549
+error rate 3.3 24.1
+LS $\alpha=0.3$samples 6726 2362
+#errors 258 647
+error rate 3.8 27.4
 Figure 6: Left: Evaluation on the combination of ImageNet + ImageNet-sketch.
@@
 During deployment the logits are normalised by their $p$-norm and then the MSP score $\pi_{\text{max}}$ is replaced by the normalised max logit,
-$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=\boldsymbol{v}/\left\lVert\boldsymbol{v}\right\rVert_{p}=\boldsymbol{v}\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
+$U=-v^{\prime}_{\text{max}}=-\max\nolimits_{k}v^{\prime}_{k},\quad\boldsymbol{v}^{\prime}=[\boldsymbol{v}+s]/\left\lVert\boldsymbol{v}+s\right\rVert_{p}=[\boldsymbol{v}+s]\big{/}\bigl{(}\sum\nolimits_{k}|v_{k}+s|^{p}\bigr{)}^{\frac{1}{p}}~{},$ (16)
 […]
-Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. A.2.
+Figure 7: RC curves with inference-time logit normalisation. Logit normalisation improves SC performance on LS models, but has little effect on CE models. For other models, see Sec. B.2.
 Figure 8: Visualisation of the effect of logit normalisation for segmentation (same scene as Fig.
@@
 Logit normalisation significantly reduces the overconfidence of the LS-trained model, although it has little effect on the uncertainty of the CE-trained model.
-We provide more figures in Appendix F.
+We provide more figures in Appendix G.
 5.1 Explaining the effectiveness of logit normalisation.
 Although Cattelan & Silva (2024) empirically validate this approach on a large number of pre-trained models, it remains unclear as to why it is so effective, or why it isn’t effective sometimes (further discussion in Sec.
-E.5).
+F.4).
 Let us shed light on the interaction between logit normalisation and LS.
@@
 In fact, we have the following inequality for positive [^6] Although this assumption may not necessarily hold, we find that in practice, $\pi_{\text{max}}$ and $v^{\prime}_{\text{max}}$ are dominated by the larger positive logits, meaning the behaviour discussed still occurs empirically (Secs.
-E.3 and 20).
+F.3 and 22).
 logits $\boldsymbol{v}$:
@@
 $v^{\prime}_{\text{max}}(\boldsymbol{v})=||\boldsymbol{v}||_{\infty}\big{/}||\boldsymbol{v}||_{p}\quad>\quad||\boldsymbol{v}+\eta\boldsymbol{1}||_{\infty}\big{/}||\boldsymbol{v}+\eta\boldsymbol{1}||_{p}=v^{\prime}_{\text{max}}(\boldsymbol{v}+\eta\boldsymbol{1})~{}.$ (18)
-The proof can be found in Appendix D.
+The proof can be found in Appendix E.
 Recalling that $||\boldsymbol{v}||_{\infty}=v_{\text{max}}$, Result 1 thus implies that for a given softmax output $\boldsymbol{\pi}(\boldsymbol{v})=\boldsymbol{\pi}(\boldsymbol{v}+\eta\boldsymbol{1})$ and arbitrary corresponding $\boldsymbol{v}$, the greater the value of $\eta$ and thus $v_{\text{max}}$, the lower the value of the normalised max logit $v^{\prime}_{\text{max}}$.
@@
 Recall how in Sec.
-4.2 we show how imbalanced logit regularisation leads to degradation in SC.
+4.2 we show how imbalanced logit suppression leads to degradation in SC.
 In particular Fig.
@@
 We see now that logit normalisation will increase the uncertainty (ranking) of errors ✗  relative to correct predictions ✓  using the information in $v_{\text{max}}$, leading to improved SC for an LS model.
-That is to say logit normalisation effectively reverses the effect of the imbalanced logit regularisation from LS, improving SC.
+That is to say logit normalisation effectively reverses the effect of the imbalanced logit suppression from LS, improving SC.
 We note that Cattelan & Silva (2024) find that models that are less confident (on both ✓ and ✗) tend to benefit more from logit normalisation, again pointing towards LS.
@@
 In the case of SC, misclassifications are to be rejected (El-Yaniv & Wiener, 2010).
-The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023).
+The baseline approach is to use the MSP (Hendrycks & Gimpel, 2017; Geifman & El-Yaniv, 2017) and there have been a number of proposed training (Moon et al., 2020; Huang et al., 2020a; Ziyin et al., 2019; Zhu et al., 2024) and architectural (Geifman & El-Yaniv, 2019; Corbière et al., 2019) enhancements, however, recently the effectiveness of some of these enhancements has been called into question (Feng et al., 2023).
 Another scenario is out-of-distribution (OOD) detection (Yang et al., 2021; Hendrycks & Gimpel, 2017), where data from outside of the training distribution are to be rejected.
@@
 Our experiments across various tasks and architectures show that LS leads to consistent degradation in a model’s ability to reject misclassifications, even if it improves accuracy.
-By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS regularises/suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions.
+By analysing the logit-level gradients of the LS loss, we provide an explanation for this previously not understood behaviour – LS suppresses the max logit more(less) the more likely a prediction will be correct(wrong), degrading a model’s uncertainty ranking of correct ✓ vs incorrect ✗ predictions.
 We then investigate post-hoc logit normalisation as a method to improve the degraded SC performance caused by LS.
-We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced regularisation by increasing the uncertainty when the max logit is higher.
+We find it to be highly effective and shed light on why – it reverses the effect of the aforementioned imbalanced suppression by increasing the uncertainty when the max logit is higher.
 We hope that our work encourages more research into understanding how different training techniques may impact model performance in downstream applications such as uncertainty estimation.
+A further discussion about the practical impact of our work, as well as potential future work can be found in Appendix H.
+Acknowledgments
+Guoxuan Xia’s PhD is funded jointly by Arm and the EPSRC.
+This work was performed using HPC resources from GENCI-IDRIS (Grant 2023-[AD011011970R3]).
+We thank the anonymous reviewers for their valuable feedback.
 Reproducibility statement
-To ensure transparency, we use publicly available datasets: CIFAR-100, ImageNet, and Cityscapes.
-Our detailed experimental methods are outlined in Appendix B.
-Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix D.
-To help replicate our work, we will soon share our source code on GitHub, including configuration files and the training code. We also plan to release our most important models on Hugging Face.
+To ensure transparency, we use publicly available datasets: CIFAR-10/100, ImageNet, Cityscapes, IMDB, Bank Marketing, and Online shoppers.
+Our detailed experimental methods are outlined in Appendix C.
+Finally, the proof supporting our explanation of the effectiveness of logit normalisation is provided in Appendix E.
+To help replicate our work, we share our source code on GitHub based on TorchUncertainty (Lafage et al., 2025), [^7] https://github.com/ENSTA-U2IS-AI/Label-smoothing-Selective-classification-Code, including the configuration files and training code.
+We also release the most important models on Hugging Face.
+[^8] https://huggingface.co/ENSTA-U2IS/Label-smoothing-Selective-classification
 Ethics
@@
 While our work aims to improve the robustness of ML systems, we acknowledge the potential risk that these advancements could be misapplied in harmful ways.
-Appendix A Additional results
-A.1 Illustrative examples of ImageNet overconfidence
+Appendix A Glossary of Notation
+We summarise the main notations used in the paper in Tab. 1.
+Table 1: Glossary of Notation
+Notation Meaning $p(\cdot)$ Probability density function $P(\cdot)$ Probability mass function $\boldsymbol{x}$ Input datum in $\mathbb{R}^{D}$ $y$ Label (categorical) from the set of $K$ possible labels $\{\omega_{k}\}_{k=1}^{K}$ $p_{\text{data}}(\cdot)$ True data distribution $\mathcal{D}_{\text{tr}}=\{\boldsymbol{x}^{(n)},y^{(n)}\}_{n=1}^{N}$ Training dataset of inputs and labels drawn from $p_{\text{data}}(\boldsymbol{x},y)$ $\boldsymbol{\theta}$ Model parameters $\boldsymbol{v}$ Logits (pre-softmax model outputs) in $\mathbb{R}^{K}$ $v_{\text{max}}$ Maximum logit, $\max\nolimits_{k}v_{k}(\boldsymbol{x};\boldsymbol{\theta})$ $\bar{\boldsymbol{\pi}}(\boldsymbol{x})$ True categorical conditional data distribution, $\bar{\pi}_{k}=P_{\text{data}}(\omega_{k}|\boldsymbol{x})$ $\boldsymbol{\pi}(\boldsymbol{x};\boldsymbol{\theta})$ Softmax output in $[0,1]^{K}$ of model, $\pi_{k}=\frac{\exp v_{k}}{\sum_{i}\exp v_{i}}=P(\omega_{k}|\boldsymbol{x};\boldsymbol{\theta})$ ()(-)( - )MSP/$\pi_{\text{max}}$ Maximum softmax probability, $\max\nolimits_{k}\pi_{k}(\boldsymbol{x};\boldsymbol{\theta})$, negate for uncertainty $\delta_{ij}$ Kronecker delta, $\delta_{ij}=1\text{ if }i=j,\text{ and }0\text{ if }i\neq j$ KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] Kullback–Leibler divergence $\mathcal{L}_{\text{CE}}(\boldsymbol{\theta})$ Empirical cross entropy loss minimised over finite data $\mathcal{D}_{\text{tr}}$ $\mathcal{L}_{\text{CE}}^{\text{true}}(\boldsymbol{\theta})$ True cross entropy loss minimised over $p_{\text{data}}(\boldsymbol{x},y)$.
+$\alpha$ Label smoothing parameter in $[0,1]$ $\mathcal{L}_{\text{LS}}(\boldsymbol{\theta};\alpha)$ Label smoothing loss $\boldsymbol{x}^{*}$ Test input datum $f(\boldsymbol{x})$ Classifier function $\hat{y}$ Label predicted by model for a given input, $\hat{y}=f(\boldsymbol{x}^{*};\boldsymbol{\theta})=\operatorname*{arg\,max}\nolimits_{\omega}P(\omega|\boldsymbol{x}^{*};\boldsymbol{\theta})$ $\bar{\pi}_{\hat{y}}$ True probability of label predicted by model $\bar{\pi}_{\hat{y}}=P_{\text{data}}(\hat{y}|\boldsymbol{x})$ $P_{\text{error}}$ Probability of a given label prediction being incorrect, $P_{\text{error}}=1-\bar{\pi}_{\hat{y}}$ $U(\boldsymbol{x})$ Scalar uncertainty score $g(\boldsymbol{x};\tau)$ Binary rejection function, 0 (reject) if $U>\tau$ otherwise 1 (accept) $\text{Risk}(f,g;\tau)$ Average error on accepted samples for threshold $\tau$ $\text{Coverage}(g;\tau)$ Proportion of all data that is accepted for threshold $\tau$ $\frac{\partial\mathcal{L}_{\text{sup}}}{\partial v_{k}}$ Suppression logit gradient, difference between LS and CE gradients, $\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}$ $\boldsymbol{v}^{\prime}$ $p$-normalised logits, $\boldsymbol{v}^{\prime}=[\boldsymbol{v}+s]/\left\lVert\boldsymbol{v}+s\right\rVert_{p}$ with scalar shift $s$.
+Appendix B Additional results
+B.1 Illustrative examples of ImageNet overconfidence
 Fig. 9 shows a few examples of overconfident misclassifications on the ImageNet evaluation data by our LS-trained ResNet-50.
@@
 3.
-A.2 Complete results for ViT and DeepLabV3+
+B.2 Complete results for ViT and DeepLabV3+
 We include additional experimental results that mirror those found in the main paper on ResNet-50, for ViT-S-16 and DeepLabV3+:
@@
 Although it is less obvious, it is clear for both ViT-S-16 and DeepLabV3+ that for higher $\pi_{\text{max}}$ the standard deviations overlap much less than for CE.
-  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix F.
-Figure 10: RC curves when evaluating on the combination of ImageNet and ImageNet-sketch.
-We see that for low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced.
+  • Fig. 13 shows the effectiveness of logit normalisation for ViT-S-16 on ImageNet and DeepLabV3+ (ResNet-101) on Cityscapes. We also provide 2 additional segmentation figures in Appendix G.
+Figure 10: RC curves evaluating on the combination of ImageNet and ImageNet-sketch.
+For low-uncertainty predictions, the degradation caused by LS is exacerbated when distribution shift is artificially introduced.
 This shows that LS leads to increasing overconfidence on less-well-fit data.
@@
 ImageNet Sketch
-CE #samples 8994 1094
-#errors 93 80
-error rate 1.0 7.3
-LS $\alpha=0.1$samples 8561 1527
-#errors 187 211
-error rate 2.2 13.8
-LS $\alpha=0.2$samples 8171 1917
-#errors 244 365
-error rate 3.0 19.0
-LS $\alpha=0.3$samples 7915 2173
-#errors 297 541
-error rate 3.8 24.9
+CE #samples 7940 1148
+#errors 94 74
+error rate 1.2 6.4
+LS $\alpha=0.1$samples 7422 1666
+#errors 174 296
+error rate 2.3 17.8
+LS $\alpha=0.2$samples 6808 2280
+#errors 225 549
+error rate 3.3 24.1
+LS $\alpha=0.3$samples 6726 2362
+#errors 258 647
+error rate 3.8 27.4
 ViT-S-16 @10% coverage of ImageNet + Sketch
 ImageNet Sketch
-CE #samples 9528 560
-#errors 88 43
-error rate 0.9 7.7
-LS $\alpha=0.1$samples 9332 756
-#errors 130 83
-error rate 1.4 11.0
-LS $\alpha=0.2$samples 9117 971
-#errors 177 164
-error rate 1.9 16.9
-LS $\alpha=0.3$samples 8966 1122
-#errors 208 195
-error rate 2.3 17.4
+CE #samples 8529 559
+#errors 90 35
+error rate 1.1 6.3
+LS $\alpha=0.1$samples 8229 859
+#errors 158 102
+error rate 1.9 11.9
+LS $\alpha=0.2$samples 7927 1161
+#errors 196 161
+error rate 2.5 13.9
+LS $\alpha=0.3$samples 7618 1470
+#errors 254 294
+error rate 3.3 20.0
 Figure 11: Distribution of the max logit $v_{\text{max}}$ given MSP $\pi_{\text{max}}$ for correct  and incorrect  predictions separately for ViT-S-16.
@@
 However, logit normalisation makes little difference to the CE-trained model.
-A.3 Small-scale CIFAR experiments
+B.3 Small-scale CIFAR experiments
 We also include experimental results on small-scale $32\times 32$ CIFAR-100 (Krizhevsky, 2009).
@@
 Figure 16: DenseNet on CIFAR-100 – LS leads to higher $v_{\text{max}}$ on errors
-Appendix B Reproducibility
-We recall that all the code used in this project will be made available soon to ensure exact reproducibility. We also plan to release our most important models on Hugging Face.
-Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix A.
-All our models are trained with PyTorch, and we use the native implementation from the CrossEntropyLoss for label smoothing.
-B.1 Image classification
+B.4 Small-scale tabular data experiments
+To increase the scope of our experiments on the degradation of selective classification due to label-smoothing, we perform small-scale tabular binary classification.
+Following the setting of TabTransformer (Huang et al., 2020b), we train two-hidden-layer MLPs on Bank Marketing (Moro et al., 2014) and Online purchasing shoppers intentions (Sakar et al., 2019).
+Figure 17: RC curves for MLPs trained on Bank Marketing and Online Shoppers using $\alpha=0.6$, hence with a gap between the positive and negative smooth labels of 0.4.
+In this binary classification setting, we set $\alpha$ to 0.6 to sufficiently reduce the gap between the optimal MSP corresponding to the positive and negative detections of the class at stake.
+With $\alpha=0.6$ the minimum of the loss with label-smoothing is achieved for a prediction with a score of 0.7 when the hard label equals 1, and a score of 0.3 when the hard label is 0, resulting in a gap of a magnitude of 0.4.
+In Fig.
+17, we see that in both cases, the AURC of the LS models are worse than those trained with the classical cross-entropy despite their accuracy being very similar.
+Furthermore, we see that the CE models are able to get much closer to zero risk, similar to the image classification experiments.
+However, we note that the effect of label smoothing seems less pronounced for small-scale tabular data.
+We posit that this may be due to the data distributions being simpler and easier to capture, such that the neural networks are generally better fit and more knowledgeable about the data.
+Thus reducing the level of imbalanced suppression over training data (see Sec.
+4).
+To address this lack of difficulty in fitting the distribution, we reduce the number of training points (as described in Appendix C).
+Moreover, training small-scale and simple datasets drastically reduces the quality of our estimation of the Risk-Coverage curves.
+Given the limited number of data and the very high accuracy achieved by our MLPs, the number of errors used to estimate the risk is limited.
+The estimation of the “true” risk-coverage curve – which would be obtained with the whole distribution – is therefore imprecise and suffers from an important variability.
+The results may thus differ when starting the optimization process from different initializations, using a different batch composition and order, or due to the non-deterministic nature of some algorithms used to compute the backpropagation (Laurent et al., 2023).
+B.5 Small-scale natural language data
+With also provide small-scale natural language processing experiments training LSTM-based models (Sepp Hochreiter, Jürgen Schmidhuber, 1997) combined with two-layer perceptrons.
+We focus on LSTMs to add another architecture, and given that CNNs and transformers were already used in the image-classification setting (see e.g.
+Sec.
+4).
+Figure 18: RC curves for LSTMs trained the IMDB Movie review dataset, using $\alpha=0.6$, hence with a gap between the positive and negative smooth labels of 0.4.
+We train two networks on the IMDB Movie Review dataset (Maas et al., 2011) with the classical binary cross-entropy loss and our implementation of the binary cross entropy with label-smoothing.
+Similarly to our results in image classification and tabular data classification, Fig.
+18 shows that the LSTMs trained with label smoothing have worse AURC than models trained with cross-entropy despite their lower error-rate.
+Furthermore, the models also display greater error rates at high-confidence (low-coverage).
+We provide more extensive details on the training in Appendix C.
+Appendix C Reproducibility
+[…]
+Here, we provide the full details of our training recipes used in the main paper and in the additional results presented in Appendix B.
+All our models are trained with PyTorch (Paszke et al., 2019).
+We use the native implementation from the CrossEntropyLoss for label smoothing in the multi-class setting and a custom version of the BCEWithLogitsLoss for the binary experiments as the original does not support label smoothing.
+We use the original implementation of the authors for the negative label smoothing (Wei et al., 2022) experiments.
+NB: The results presented in the final camera-ready ICLR 2025 version of this paper may differ slightly to earlier preprint versions.
+This is due to models being retrained with different random seeds, train-val-test splits, and having minor recipe and hardware differences after a consolidation of the experimental codebase to TorchUncertainty.
+[^11] https://torch-uncertainty.github.io/ Readers should use the information in this final version for reproducibility purposes.
+C.1 Image classification
 DenseNet – CIFAR100.
-For the DenseNet trained on CIFAR-100, we extract a subset of CIFAR’s training set containing 5000 images to create a proper validation set.
+For the DenseNet trained on CIFAR-100, we randomly split the original test set into 2000 images for validation and 8000 for evaluation.
 We take batches of 64 32$\times$32-pixel images and train on a single GPU for 300 epochs using stochastic gradient descent and a starting learning rate of 0.1, Nesterov (Nesterov, 1983; Sutskever et al., 2013), a momentum of 0.9, and $1\times 10^{-4}$ weight decay.
 We divide the learning rate by ten after 150 and 225 epochs.
-We use standard augmentations.
-We apply random crop with a four-pixel padding as well as random horizontal flip.
+We use standard augmentations: we apply random crop with a four-pixel padding as well as random horizontal flip.
 We do not perform model selection and keep the last checkpoint.
 ViT-S-16 – ImageNet.
-For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 8 V100 for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999.
+For our ViT-S-16 trained on ImageNet, we take batches of 2048 images and train on 4 A100s for 300 epochs with AdamW (Loshchilov & Hutter, 2019) with the $\beta$s equal to 0.9 and 0.999.
 We start with a linear warmup for 15 epochs, then use a cosine annealing scheduler with $3\times 10^{-3}$ as the starting learning rate.
@@
 For ImageNet, we do not perform model selection and keep the last checkpoint.
-However, we randomly extract a validation set of 50000 images from the training set to perform logit normalisation.
+However, we randomly extract a validation set of 10,000 images from the validation set to perform logit normalisation, and evaluation on the remaining 40,000.
 ResNet-50 – ImageNet.
-Our ResNet-50 is trained on 8 V100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images.
+Our ResNet-50 is trained on 4 A100 with stochastic gradient descent for 120 epochs using a batch size of 1024 images.
 After five epochs of linear warmup, we use a cosine annealing scheduler starting with a learning rate of 0.4 with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$.
@@
 We use the same validation set as for the ViT for logit normalisation.
-B.2 Semantic segmentation
+C.2 Semantic segmentation
 Deeplabv3+.
@@
 The weights are optimised with a momentum of 0.9 and a weight decay of $10^{-4}$.
-We take a batch size of 12 images and train for 40000 steps.
+We take a batch size of 16 images and train for 30,000 steps.
 During training, we randomly crop the input images and targets to squares with 768-pixel-long sides.
@@
 For testing, we use the images at their original resolution and do not perform any test time augmentations.
-For the RC curves, we randomly sample 5000 predictions per image, extracted prior to the final interpolation to reduce correlations between the predictions.
+For the RC curves, we randomly sample 5000 predictions per image extracted prior to the final interpolation to compute the coverage and error rates.
 We keep the pixel-wise locations of the samples when changing the level of label smoothing $\alpha$ to ensure fair comparisons.
-B.3 Logit Normalisation
+C.3 Tabular data
+We perform the tabular data binary classification on two UCI datasets: Bank Marketing and Online shoppers.
+These two datasets have an input dimension and a number of samples of 7 and 45,211 for the former and 16 and 12,330 for the latter.
+In both cases, we select 80% of the data points for the test set and train the models for 10 epochs using PyTorch’s default Adam optimizer and a batch size of 128.
+Similarly to the other experiments of this paper, we do not perform model selection and keep the last checkpoint.
+Our models are MLPs with two hidden layers, whose size depends on the input size as follows: the first hidden layer has 4 times as many neurons as the dimension of the input data, and the second has half the number of neurons of the first layer.
+C.4 Natural language data
+For the training on the IMDB (Maas et al., 2011) dataset, we tokenize the data with nltk and convert them to 300-dimensional vectors using GloVe 840B (Pennington et al., 2014).
+The architecture of the models is defined as follows:
+  • an LSTM layer with a 300 input dimension and a 256 output dimension, followed by a dropout layer of rate 0.2,
+  • a linear layer keeping the dimension of the vectors unchanged, on which we apply ReLU, followed by a second dropout layer with the same rate,
+  • two linear layers with ReLU reducing the dimension to 128 and 1, respectively (binary classification).
+Finally, we train the models with PyTorch’s Adam default for 10 epochs using the pre-made train test split. The code is available on GitHub.
+C.5 Logit Calculations
+We perform logit evaluation calculations in double precision (softmax, logit normalisation, entropy etc.), in order to reduce the impact of numerical error.
 […]
-Appendix C Analysis of empirical loss (one-hot labels)
+Appendix D Analysis of empirical loss (one-hot labels)
 We can perform a similar analysis as in Sec.
@@
 $\frac{\partial\mathcal{L}_{\text{CE}}}{\partial v_{k}}=-\left[\delta_{y\omega_{k}}\!\!-\pi_{k}\right],\quad\frac{\partial\mathcal{L}_{\text{LS}}}{\partial v_{k}}=-\biggl{[}\Bigl{[}\,\underbrace{(1-\alpha)\delta_{y\omega_{k}}}_{\text{{\color[rgb]{0.58203125,0.41796875,0.9296875}data supervision}}}+\!\!\!\!\underbrace{\alpha/K}_{\text{{\color[rgb]{0.61328125,0.5,0.0234375}regularisation}}}\!\!\!\!\Bigr{]}\!-\pi_{k}\biggr{]},$ (19)
-which in turn gives the regularisation gradient,
-$\frac{\partial\mathcal{L}_{\text{reg}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
+which in turn gives the suppression gradient,
+$\frac{\partial\mathcal{L}_{\text{sup}}}{\partial v_{k}}=\frac{\partial(\mathcal{L}_{\text{LS}}-\mathcal{L}_{\text{CE}})}{\partial v_{k}}=\frac{\partial\mathcal{L}_{\text{LS}}}{\partial v_{k}}-\frac{\partial\mathcal{L}_{\text{CE}}}{\partial v_{k}}=\alpha\left[\delta_{y\omega_{k}}-1/K\right]~{},$ (20)
 which is once again independent of the model output $\boldsymbol{\pi}$.
-In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different regularisation gradients for the max logit $v_{\text{max}}$,
+In this case, as the label $y$ has already been sampled, the model is either right  or wrong , giving two different suppression gradients for the max logit $v_{\text{max}}$,
 […]
-We can see that the max logit is more strongly regularised/suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper.
+We can see that the max logit is more strongly suppressed during training when the prediction is correct, which aligns with the analysis of $\mathcal{L}^{\text{true}}$ and $P_{\text{error}}$ in the main paper.
 When the model is correct during training, the max logit is suppressed, but when it is incorrect the max logit is not suppressed.
@@
 5 (right) where $v_{\text{max}}$ is higher for errors given the value of MSP.
-Appendix D Result and proof
+Appendix E Result and proof
 Result.
@@
 $\frac{(v_{k}+\eta)^{p}}{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}\geq\frac{v_{k}^{p}}{\max\limits_{k\leq K}v_{k}^{p}}.$ (25)
-As for Eq. 23, please note that using $k=i$, we get the same equation as Eq. 25, although with a strict inequality. We can now sum on the elements of $\boldsymbol{v}$ to get
+Similar to Eq. 23, please note that using $k=i$, we get the same equation as Eq. 25, although with a strict inequality. We can now sum on the elements of $\boldsymbol{v}$ to get
 $\frac{\sum\limits_{k=1}^{K}(v_{k}+\eta)^{p}}{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}>\frac{\sum\limits_{k=1}^{n}v_{k}^{p}}{\max\limits_{k\leq K}v_{k}^{p}}.$ (26)
-By taking the inverse, we get
+By taking the inverse (all values are strictly positive), we get
 $\frac{\max\limits_{k\leq K}v_{k}^{p}}{\sum\limits_{k=1}^{n}v_{k}^{p}}>\frac{\max\limits_{k\leq K}(v_{k}+\eta)^{p}}{\sum\limits_{k=1}^{K}(v_{k}+\eta)^{p}}.$ (27)
@@
 ∎
-Appendix E Discussions
-E.1 $U$ other than MSP
+Appendix F Discussions
+F.1 $U$ other than MSP
 In the main body of the paper we focus solely on MSP as our uncertainty score $U$.
@@
 Finally, we also choose to omit various training/architecture-based approaches such as (Ziyin et al., 2019; Geifman & El-Yaniv, 2019; Moon et al., 2020; Zhu et al., 2024), as we aim to focus our investigation solely on understanding the training effects of LS.
-We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use-cases and so may be more likely to be chosen by a practitioner.
+We note that LS (potentially combined with post-hoc logit normalisation) is considerably simpler to implement than the aforementioned training-based approaches, and has been shown to be stable in many more use cases and so may be more likely to be chosen by a practitioner.
 Besides, recent work (Feng et al., 2023) has suggested that MSP applied to the methods in (Ziyin et al., 2019; Geifman & El-Yaniv, 2019) actually performs better than the $U$s proposed in them (reject logit/selection head).
@@
 This behaviour is in line with existing work that shows that uncertainty scores designed for OOD detection are not suitable for SC.
-E.2 Aleatoric and epistemic uncertainty
+F.2 Aleatoric and epistemic uncertainty
 We recognise that the conceptual decomposition of predictive uncertainty into aleatoric and epistemic uncertainty (Gal, 2016; Hüllermeier & Waegeman, 2021; Kirsch, 2024) can be applied to the discussion throughout this paper, and that some readers may be confused as to why we do not use these terms.
 We choose to use simpler, more direct language as we believe it more efficiently conveys our discussion, reduces the number of concepts to introduce and also reduces any potential confusion.
-E.3 On the assumption that logits are positive
+F.3 On the assumption that logits are positive
 Figure 21: Mean±$\pm$±std after logits have been sorted highest to lowest for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set.
@@
 of $v^{\prime}_{\text{max}}$ in a 0.2-wide sliding window for samples with $v_{\text{max}}$ within mean±$\pm$2std of $v_{\text{max}}$ within the bin (to remove noisy averages).
+Largest positive logits dominate.
 Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true.
@@
 This is shown in Fig.
-19, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case).
-Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive.
-As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
+21, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case).
+Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive and dominate the computation.
+As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$, even if not all $v_{k}>0$.
+(Note that we omit shift $s=-1/K\sum_{k}v_{k}$ as it is $\approx 0$ in our experiments).
 We can also consider the scenario where we add $\eta$ to only the top-$k$ logits.
 This more aptly describes the empirical logit behaviour compared to adding $\eta$ to all logits, as the values of the lower ranking logits vary much less than the higher ranking ones (Fig.
-19).
+21).
 Here we would expect $\pi_{\text{max}}(\boldsymbol{v})$ to increase very slightly, but would expect $v^{\prime}_{\text{max}}(\boldsymbol{v})$ to decrease as the numerator of Eq.
@@
 Thus, logit normalisation is able to improve the SC performance of LS-trained models by penalising the confidence of higher max logit values.
-E.4 Existing benchmarks and training recipes with LS
+Logit centralisation.
+We follow Cattelan & Silva (2024) and suggest to shift the logits by their mean during normalisation,
+$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k}.$ (28)
+In our experiments, this has little-to-no effect as $\mu\approx 0$ (for example for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation data $\mu$ averages $\sim 3\times 10^{-3}$ with std.
+$\sim 1\times 10^{-3}$).
+This is corroborated by Cattelan & Silva (2024)’s observations in their Appendix G, where they find this shift does not affect the vast majority of their models.
+We note that according to the analysis in Sec.
+5.1 and Sec.
+F.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits.
+Logits are not necessarily constrained to be positive.
+They can be arbitrarily shifted without affecting the cross entropy loss (e.g.
+via bias $\boldsymbol{b}$ of the final layer), or derived as log probabilities which are necessarily negative.
+Thus we suggest applying the above shift to ensure that the largest logits are positive.
+We note that other shifts are possible (for example $s=-\min_{k}v_{k}$) to achieve the same purpose but we leave the exploration of this to future work.
+F.4 Explanations of Logit normalisation in Cattelan & Silva (2024)
+The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways.
+However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B.
+They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix E).
+We believe their explanation is ultimately incomplete as:
+  1. 1.
+    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification.
+This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties.
+It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples (Zhu et al., 2024).
+  2. 2.
+    Although they elucidate some of the mechanics of logit normalisation, showing that logit normalisation reduces the impact of smaller logits on uncertainty, they do not link it to any observed model behaviour that differs between correct ✓ vs incorrect ✗ predictions, instead just speculating that models may be underconfident on certain sets of samples.
+To explain why an approach improves SC, we need to know how it treats ✓ and ✗ differently to improve the rank ordering of uncertainties between ✓ and ✗.
+On the other hand, our explanation focuses solely on the relative ranking of uncertainties between ✓ vs ✗ samples and is able to link the mechanics of logit normalisation to how the behaviour of correct ✓ vs incorrect ✗ predictions differ under label smoothing (Figs.
+5 and 1).
+F.5 Analysis of Negative Label Smoothing (Wei et al., 2022)
+We consider the framework developed in the main paper to study Generalized Label Smoothing (GLS) with negative smoothing values – called Negative Label Smoothing (NLS) – as suggested by Wei et al.
+(2022).
+The definition of GLS is the same as the original from Szegedy et al.
+(2016), except for the domain of the label smoothing value $\alpha$, which can take values in $(-\infty,1]$.
+Wei et al.
+(2022) suggest that using negative values for the label smoothing value can lead to improved performance when training with noisy labels as well as in the “clean” setting.
+It is tempting to consider, as negative $\alpha$ ought to reverse the logit suppression described in Eq.
+15.
+However, here we show that NLS can lead to unstable training characteristics.
+The per-sample cross-entropy between a single softmax prediction $\pi$ and the corresponding true categorical distribution $\bar{\pi}$ is convex and can be written as follows:
+$\text{CE}(\pi,\bar{\pi})=-\sum_{k\in\llbracket 1,K\rrbracket}\left[(1-\alpha)\bar{\pi}_{k}+\frac{\alpha}{K}\boldsymbol{1}\right]\log\pi_{k}.$ (29)
+For $\alpha\geq 0$, which covers vanilla cross entropy ($\alpha=0$) and regular label smoothing ($\alpha>0$) a single global minimum is attained in $\pi=(1-\alpha)\bar{\pi}+\frac{\alpha}{K}\boldsymbol{1}$.
+Here, the gradient of the loss is zero therefore creating a balance and stabilizing training, taking the gradient as in Eq.
+13,
+${\color[rgb]{0,0,0}\frac{\partial\text{CE}(\pi,\bar{\pi})}{\partial v_{k}}=-\biggl{[}\underbrace{\Bigl{[}(1-\alpha)\bar{\pi}_{k}+\alpha/K\Bigr{]}}_{\text{target}}\!-\underbrace{\pi_{k}}_{\text{softmax output}}\biggr{]}.}$ (30)
+However, in the case of a negative label smoothing value $\alpha<0$ the target can be outside $[0,1]$, but the softmax output is constrained to $[0,1]$, thus $\pi=(1-\alpha)\bar{\pi}+\frac{\alpha}{K}\boldsymbol{1}$ is not achievable and the gradient will always have some magnitude and there is no optimisation minimum.
+Indeed, if we consider for one-hot $\bar{\pi}$, the $i$-th terms s.t.
+$\bar{\pi}_{i}=0$ of Eq.
+29 will be strictly negative with a negative-label-smoothed target.
+Since the log of the softmax probabilities $\pi$ can take values in $(-\infty,0]$, so can the loss.
+Being convex, there cannot exist a local minimum, and therefore, the gradient can never be zero and pushes its value to $\infty$- ∞ by reducing the value of the corresponding logit to $\infty$- ∞ and pushing the others to $\infty$.
+We hypothesise that this unbalance explains the instability of the method that users reported on the GitHub repository of the original paper.
+Unfortunately, we too were not able to reproduce the experiments of Wei et al.
+(2022) and obtain stable enough training runs to perform experiments on NLS and confirm this theoretical analysis.
+F.6 Existing benchmarks and training recipes with LS
 Although we do not exhaustively search all training recipes for all models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024), we do provide a number of examples of evaluated models trained with label smoothing.
@@
 However, after inspecting both the original papers and open-source repositories [^12] https://github.com/google-research/vision_transformer of the aforementioned work we were unable to find any confirmation of the use of label smoothing.
-E.5 Logit normalisation in (Cattelan & Silva, 2024)
-Explanations of effectiveness.
-The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways.
-However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B.
-They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix D).
-We believe their explanation is ultimately incomplete as:
-  1. 1.
-    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification.
-This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties.
-It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples Zhu et al.
-(2024).
-  2. 2.
-    Although they elucidate some of the mechanics of logit normalisation, showing that logit normalisation reduces the impact of smaller logits on uncertainty, they do not link it to any observed model behaviour that differs between correct ✓ vs incorrect ✗ predictions, instead just speculating that models may be underconfident on certain sets of samples.
-To explain why an approach improves SC, we need to know how it treats ✓ and ✗ differently to improve the rank ordering of uncertainties between ✓ and ✗.
-On the other hand, our explanation focuses solely on the relative ranking of uncertainties between ✓ vs ✗ samples and is able to link the mechanics of logit normalisation to how the behaviour of correct ✓ vs incorrect ✗ predictions differ under label smoothing (Figs.
-5 and 1).
-Logit centralisation.
-Cattelan & Silva (2024) suggest an additional step of logit centralisation,
-$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k},$ (28)
-which in practice brings most of the lower valued logits close to zero if they are not already.
-They note that for the majority of models this step is unnecessary, and we also find it unnecessary in our experiments so we omit it for the sake of simplicity.
-We note that according to the analysis in Sec.
-5.1 and Sec.
-E.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits, which centralisation seems to empirically ensure by making most of the smallest logits close to zero (Cattelan & Silva, 2024).
-Appendix F Additional segmentation results
-In this section, we complete the picture by providing additional semantic segmentation results to Figures 1 and 8.
-As in the main paper, the following predictions are performed by a ResNet-101-based DeepLab-v3+ trained on Cityscapes.
+Appendix G Additional segmentation results
+In this section, we complete the picture by providing additional semantic segmentation results to Figs.
+1 and 8.
+As in the main paper, the following segmentation maps are performed by ResNet-101-based DeepLab-v3+ trained on Cityscapes with cross-entropy and label-smoothing with $\alpha=0.2$.
+We recall that the models and notebooks used to generate these plots are available on Hugging Face and GitHub, respectively.
 In these figures, we provide the rank of all the predictions – not only errors – to show the difference between the model’s behavior on its errors but also correct predictions.
 The rank of the correct predictions ✓ is slightly greyed out compared to the rank of the errors ✗.
-Figure 21 presents the full ranks corresponding to the scene used in the main paper.
-We will provide other examples to the readers upon request.
-Later, we will make our models available to ease the reproducibility of our work.
+Fig.
+23 presents the full ranks corresponding to the scene used in the main paper.
 Figure 23: We provide all the predicted ranks on our segmentation plots to understand where the model predicted higher relative confidence.
 The pixel uncertainty rankings provided by the model trained with CE (left) are smoother than those of the model trained with LS $\alpha=0.2$ (right).
-Figure 22 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are smoother than those of the label-smoothing-based one.
-The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly decreases nearby the boundary of objects.
+Figs.
+23 and 24 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are generally smoother than those provided by the label-smoothing-based model.
+The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly increases near the boundary of objects.
 However, the boundaries are sometimes wrongly predicted, leading to (rank-wise) highly confident errors ✗.
-The pixels corresponding to the boundaries of objects, where naturally the probability of error is higher and which are most sensitive to label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model.
+The pixels corresponding to the boundaries of objects, where the probability of error is naturally higher and which are most sensitive to (ground-truth) label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model.
 This explains the high-confidence errors ✗ in the misclassified boundaries of several objects, either in Figs.
-1 and 22, and aligns with the imbalanced regularisation of Eq.
+1 and 24, and aligns with the imbalanced suppression of Eq.
 15.
-We also provide Figure 23 to show that the logit normalisation method is not always successful in solving the ranking of the predictions.
-In this figure, we see that a large part of the misclassified pixels (lower left) keeps low uncertainty ranks either using the CE-based model or the LS-based model while using logit normalisation.
-However, there are some improvements on the label-smoothing side, where the original low-ranked boundary errors ✗ are less (rankwise) confident, such as the roof of the car and the right border of the median strip.
-Interestingly, LS-based high confidence errors in Figure 23 mainly correspond to a change of texture of the median strip – likely associated with an uncertainty on the true label – and correlate with those of the CE-based model.
-For brevity, we provide an image for a logit normalisation with norm dimension $p=5$, but the results do not differ much from the other values.
-As stated above, we will provide additional figures to readers upon request.
-Figure 22: Logit normalisation improves on boundary-related very confident errors ✗, but not on the bottom right errors ✗.
-Figure 23: Logit normalisation does not always fix selective classification when used on a model trained with a label-smoothing (right).
-However, boundary errors ✗ have lower (ranked) confidence, such as for the car’s roof.
+We again see how logit normalisation is able to decrease the confidence of the label-smoothing-trained model on highly confident (bright yellow) errors ✗.
+[…]
+Figure 24: Logit normalisation improves on boundary-related very confident errors ✗, but does not completely fix the high (ranked) confidence of the bottom right errors ✗.
+Figure 25: Logit normalisation does not always completely fix selective classification when used on a model trained with a label-smoothing (right).
+Appendix H Impact and Future Work
+In this section, we provide some additional discussion about the (practical) impact of our work as well as promising directions for potential future research.
+Direct impact.
+By empirically verifying and analytically elucidating the limited experimental results pertaining to LS in (Zhu et al., 2024) (it is not the focus of that work), we provide strong evidence that the behaviour that LS degrades SC is generalisable (over architectures, data modalities etc.).
+In particular, we directly analyse the loss, which is common to all settings involving label smoothing.
+This will help inform practitioners of selective classification when they are designing and deploying systems.
+Furthermore, by explaining the efficacy of logit normalisation we provide an effective and well-motivated solution to the previously demonstrated problem.
+We emphasise that when logit normalisation was introduced in (Cattelan & Silva, 2024), it was not clearly explained why logit normalisation was effective on some pretrained models and ineffective on others.
+In our work, we analytically clarify the mechanism of logit normalisation, opening up the figurative black box, and are able to directly link it to our previous analysis on LS.
+This provides clear guidance on when and why to use logit-normalisation, giving potential practitioners confidence in the effectiveness of the approach, which is especially important in high-risk safety-critical applications.
+Future work.
+The empirical and analytical results relating to LS in our work naturally suggest that other training approaches that alter the labels such as Mixup (Zhang et al., 2018) may also have similar adverse effects and/or be amenable to similar gradient analysis.
+It also raises the question of how such label augmentations effect problem settings outside of selective classification such as OOD detection or transfer learning.
+The novel analysis of LS in Sec.
+4 may also be useful in problem settings beyond SC: Does this aspect of label smoothing (suppressing the max logit less for incorrect predictions) affect generalisation?
+What about behaviour on OOD data?
+Could it help explain the behaviour in (Kornblith et al., 2021) where LS results in worse transfer/representation learning?
+Can this insight lead to a modification of LS to improve it?
+Our analysis of the mechanism of logit normalisation in Sec.
+5.1 is also general – we simply demonstrate that logit normalisation reduces confidence when the max logit is higher.
+Thus, this knowledge can be potentially applied to other scenarios (e.g.
+OOD detection).
+Furthermore, in the case of post-hoc methods, logit normalisation using the $p$-norm is just one possible option for correcting the effect of LS – armed with the insight presented in our work it may be possible to develop superior or more principled methods for extracting better uncertainty estimates from logits.